### v2

# Beyond Per-Querier Budgets: Rigorous and Resilient Global Privacy Enforcement for the W3C Attribution API

Abstract.
We analyze the privacy guarantees of the Attribution API, an upcoming W3C standard for privacy-preserving advertising measurement. Its central guarantee—separate individual differential privacy (IDP) budgets per querier—proves unsound once data adaptivity across queriers is considered, a condition we argue is unavoidable in practice. The issue lies not with IDP or its device-epoch unit, but with the per-querier enforcement model, which has also appeared in other DP systems; we show formally that no per-querier accounting scheme, under either individual or traditional DP, remains sound under adaptivity, a gap missed by prior analyses. By contrast, a global device-epoch IDP guarantee remains sound, and we introduce Big Bird, a privacy budget manager for the Attribution API that enforces this guarantee. The challenge is that a global budget shared across many untrusted queriers creates denial-of-service (DoS) risks, undermining utility. Building on prior work that treats global budgets as a computing resource, we adapt resource isolation and scheduling techniques to the constraints of IDP, embedding DoS resilience into the budget management layer. Our Rust implementation with Firefox integration, evaluated on real-world ad data, shows that Big Bird supports benign workloads while mitigating DoS risks. Still, achieving both utility and robustness requires global budgets to be configured more loosely than per-site budgets; we therefore recommend that the Attribution API continue using tight per-site budgets but clarify their limited formal meaning, and complement them with global budgets tuned for benign load with added slack for DoS resilience.
Contact authors: Pierre Tholoniat, Roxana Geambasu, Mathias Lécuyer ({pierre,roxana}@cs.columbia.edu, mathias.lecuyer@ubc.ca)
conference: SOSP’25; October 2025; Seoul, Republic of Korea
1  Introduction
An important new standard is taking shape within the W3C: the Attribution API (https://w3c.github.io/attribution/), designed to enable privacy-preserving advertising and other workloads that measure cross-site user activity without relying on cross-site tracking. Historically, browsers have lacked structured support for cross-site tasks such as conversion attribution measurement in advertising, which links ads viewed on content sites to purchases on seller sites, enabling evaluation and optimization of ad campaigns and placement. In the absence of such support, the advertising ecosystem has depended on invasive practices – third-party cookies, fingerprinting, and other cross-site tracking mechanisms – that undermine user privacy.
The Attribution API aims to provide a structured alternative: content sites register ads with the browser; seller sites query the browser by requesting encrypted reports of which ads a user previously saw; and reports are made available only through differentially private (DP) aggregation across many users, using secure multiparty computation or trusted execution environments. Each user’s browser controls the flow of information by deducting privacy loss from a locally maintained privacy budget before releasing an encrypted report to a querier site. Once the budget for a site is exhausted, the browser issues only encrypted null reports, thereby preventing further data flows to that site.
The theoretical foundation for this browser-controlled privacy budgeting, and the algorithm for efficiently accounting for privacy loss in-browser, was established in the Cookie Monster paper (TKM+24, 41). Cookie Monster’s key insight was to formulate guarantees in terms of individual differential privacy (IDP), a variant of DP that allows each participant—such as a browser—to maintain its own privacy budget. Because IDP measures privacy loss based on the specific contribution of a participant’s data, rather than worst-case contribution across all participants, it enables tighter accounting and better utility. For conversion attribution, where users’ contributions vary widely (different ads seen, different purchases made, and most users not participating in most measurements because they do not purchase most advertised products), this optimization was especially attractive to industry, promising to yield a useful API. Cookie Monster’s IDP-based formulation was thus rapidly adopted as the foundation of the Attribution API, which is now undergoing W3C standardization with involvement from all major browser vendors and is expected to move to public review in early 2026. If successful, the API could be deployed on billions of devices and serve as a blueprint for privacy-preserving federated measurement more broadly.
Because of the API’s importance as a standard, not just a single design or implementation, it is crucial to scrutinize its privacy guarantees. This paper makes two contributions.
As our first contribution 1 (§3), we analyze the Attribution API’s primary guarantee: a device-epoch-level [^1] In this paper, a device is a user’s browser, and an epoch is a span of time (e.g. a week) including the user’s impressions and conversions from that period. IDP guarantee enforced separately for each querier. The problem is not with IDP itself, nor with the device-epoch privacy unit, but with the per-querier enforcement model. This model is already known to be vulnerable to querier collusion, a limitation acknowledged by the API’s editors and accepted as a tradeoff. Our analysis uncovers two additional flaws, previously unknown and orthogonal to collusion, each with implications well beyond Attribution. First, when data adaptivity across queriers is possible—a condition intrinsic to streaming settings such as advertising—per-querier accounting fails, whether based on individual or traditional DP, a gap overlooked in prior analyses due to inaccurate system modeling. Second, even if adaptivity is assumed away, introducing shared limits on privacy loss spanning multiple queriers—as proposed in several DP systems (GHK+18, 14, 43, 24, 32), including Attribution, to counter collusion—is itself sufficient to invalidate the per-querier guarantees. Taken together, these results show that beyond collusion, per-querier enforcement is fundamentally unsound: data adaptivity breaks it in streaming settings, and shared limits break it even in static ones.
As our second contribution (§4), we demonstrate that a global (DP or IDP) guarantee, applied across all queriers, remains sound under these conditions and can serve as a viable option for Attribution. However, a new challenge arises: a single global budget, shared among many untrusted queriers, is vulnerable to denial-of-service (DoS) attacks that exhaust privacy resources and block benign measurement. To address this, we introduce Big Bird, a global budget management system. Big Bird achieves DoS resilience thanks to a quota system that assigns the top-level site of each conversion and impression a renewable quota drawn from the global budget. We articulate and configure the different quotas in the system to prevent attacks without being overly restrictive, by identifying structural properties of API use and a novel stock-and-flow pattern of privacy loss. This prevents monopolization while ensuring that total privacy loss never exceeds the global bound on any browser, making global IDP enforcement both sound and practical.
We implement Big Bird in two components: (1) pdslib, a generic in-browser IDP library that subsumes Cookie Monster and extends it with Big Bird ’s budget management, and (2) integration into Mozilla Firefox’s Attribution prototype. Together, these can serve as reference implementations for privacy-preserving attribution. Our evaluation, using a real-world dataset from the Criteo ad-tech company, demonstrates that: (1) well-chosen quotas preserve utility for benign workloads, (2) quotas isolate benign sites under attack, and (3) batched scheduling boosts utilization without sacrificing isolation.
Still, achieving both utility and resilience requires global budgets to be configured more loosely than baseline per-querier budgets. We therefore recommend that Attribution continue to use tight per-querier budgets, but clarify that these function as engineering constraints rather than formal privacy guarantees. Global budgets should then complement them, carrying the true formal guarantees while being set to support benign workloads and some additional slack to preserve robustness against DoS attacks.
We release our source code via several repositories: pdslib at https://github.com/columbia/pdslib, Firefox integration at https://github.com/columbia/pdslib-firefox, and experimental infrastructure at https://github.com/columbia/bigbird.
2  Background on Attribution API
Fig. 1. Attribution architecture.
2.1  Architecture
The W3C Attribution API is a browser-based interface for cross-site conversion measurement that replaces ad-tech tracking with structured, privacy-preserving mechanisms. Figure 1 illustrates the workflow involving four principals: (1) Impression sites (content providers such as news.ex or blog.ex) register ads via saveImpression(); (2) Conversion sites (advertisers such as shoes.ex) call measureConversion() to request encrypted reports of which ads a user previously saw; (3) Intermediaries (ad-techs such as adtech.ex) receive those encrypted reports and perform conversion attribution measurement; and (4) Aggregation services (e.g., divviup.org) use secure multiparty computation (MPC) or trusted execution environments (TEEs) to aggregate reports and apply differential privacy (DP) noise.
The browser mediates information flows: impressions are stored locally, conversions trigger encrypted reports, and aggregators release only DP-protected statistics. A typical sequence is: 1- 2 impression sites record ads; 3 a conversion site requests measurement via measureConversion(); 4 the browser finds relevant impressions, checks local privacy budgets, and 5 generates encrypted reports that attribute the value of the conversion to the relevant impressions; 6 reports from multiple user devices are batched, 7 submitted for aggregation, and 8- 9 released with DP noise to the intermediary and ultimately to the conversion site.
Privacy enforcement happens on-device, within the call to measureConversion(). Before any report is released, the browser deducts the corresponding privacy loss from a local privacy budget store and returns null reports when budget is exhausted. This per-device enforcement aligns with the browser’s role as the user’s privacy boundary.
2.2  From Cookie Monster to Attribution API
The Attribution API builds on the Cookie Monster framework (TKM+24, 41), which proposed leveraging individual differential privacy (IDP) (ESS15, 11, 13) as a natural fit for on-device accounting. Under IDP, each participant – here, a browser – maintains its own privacy budgets, and privacy loss is measured with respect to that individual’s actual contribution to a measurement, rather than a population-wide worst case.
This distinction proved decisive for adoption. Because users contribute unevenly to advertising data – some see many ads but never convert, some purchase products without seeing ads for them – IDP allows tighter accounting and far better utility than traditional DP. Cookie Monster provided a lightweight algorithm for efficient, IDP-based, in-browser accounting: impression and conversion data is split into non-overlapping epochs (such as one week), and privacy loss is charged only in epochs containing impressions relevant for a conversion. In those epochs, the charge is bounded by $\text{(value / maxValue)}\times\epsilon$, taking arguments of the API call such as the conversion value and DP parameters.
These properties promised a practical implementation of the API, leading the Attribution API editors to adopt Cookie Monster’s formulation as the foundation for their standardization effort. All major browsers now participate in this process, with the specification expected to be transitioned to public review in early 2026.
2.3  Queriers and privacy guarantees
Attribution performs privacy accounting with respect to a querier – the entity on whose behalf aggregation ultimately occurs. In today’s API, the querier is the advertiser (conversion site) measuring its own campaign effectiveness. A planned extension will also support ad-techs as queriers for cross-advertiser queries, such as model training for ad placement.
Each browser therefore maintains per-epoch budgets per querier, tracked in a Privacy Budget Store. Cookie Monster showed that, when modeled in isolation, such budgets implemented via individual privacy filters (FZ21, 13) satisfy device–epoch–level IDP. Google’s ARA design follows the same logic at an event-level IDP granularity (GHH+25, 16).
The Attribution specification acknowledges a known limitation of per-querier guarantees: when multiple queriers collude, their privacy losses add. To mitigate this, Attribution also defines global safety limits – per-epoch budgets spanning all queriers – as a defense-in-depth measure. The design thus treats per-querier budgets as the main engineering mechanism for controlling data flow, while relying on global limits to cap overall loss. Similar multi-granularity patterns appear in other DP systems (GHK+18, 14, 43, 24, 32).
2.4  This paper’s perspective
We show that the limitations of per-querier guarantees extend well beyond collusion. First, when data adaptivity across queriers is possible – as is inherent in advertising and other streaming workloads – no per-querier accounting scheme, under either IDP or traditional DP, remains sound. Second, even without adaptivity, adding global limits that span multiple queriers can itself undermine the validity of per-querier guarantees.
These findings expose a gap in prior analyses with implications beyond Attribution. They motivate our two contributions: (1) a formal analysis of per-querier guarantees in a general setting (next section), and (2) a practical design for rigorous and robust management of the global safety limits tailored to Attribution (§4).
3  Contribution 1: Per-Querier DP Analysis
A number of works have claimed, or at least suggested, that it is possible and valuable to enforce DP guarantees on a per-querier basis. Some offered formal backing, while others made only informal claims. Examples include Cookie Monster (TKM+24, 41), which formally analyzed per-querier IDP in Attribution; Google’s ARA paper (GHH+25, 16), which proved similar per-querier properties for an event-level variant; and systems such as DProvDB (DProvDB, 43), DPolicy (KVLH25, 24), and PSI (GHK+18, 14), which argued for multi-granularity or per-analyst enforcement in static traditional-DP database settings. These ideas make practical sense, especially in contexts where queriers are assumed to follow rules and not deliberately collude — for example, scientists querying a DP medical database under IRB protocols. In such settings, per-querier (or per-group-of-querier) guarantees appear attractive and intuitively sound.
Our analysis shows, however, that the problems with per-querier guarantees run deeper than collusion. To perform a DP analysis, one specifies: (1) the privacy property believed to hold (e.g., traditional DP, IDP, or other definition; the unit of protection, and the enforcement scope), (2) a formal model of the system’s behavior, and (3) a proof connecting the model to the property. The modeling step typically follows a particular pattern that supports analysis: queriers are treated as adversaries, a challenge bit encodes the presence or absence of a user’s data, and the model specifies the data generation and querying protocol. If the model omits essential system behaviors, however, the resulting proofs may be correct yet not reflective of the actual system.
We identify two modeling oversights in the literature on per-querier enforcement. First, analyses typically model each querier in isolation, overlooking how, in streaming settings, new data may be influenced by results from prior DP queries, sometimes across queriers. For example, if an ad placement optimized by querier $q1$ leads a user to purchase $q1$’s product, that purchase may displace the user’s purchase of another querier $q2$’s product, altering the data seen by $q2$. Second, prior work modeled per-querier budgets and global limits separately, ignoring their subtle interactions. Correcting these oversights, we show that per-querier guarantees are unsound even without collusion: adaptive data generation breaks them in one way, and shared limits break them in another. The implication is not that per-querier guarantees are useless — they remain valuable engineering constraints — but that their formal property requires stronger assumptions than previously understood. More broadly, our findings highlight the need for greater rigor in system modeling for DP, a step long recognized in formal methods but still under-appreciated in DP research.
In the remainder of this section we develop and analyze a generic model of multi-querier systems with per-querier enforcement. The model captures both adaptive data generation and shared privacy bugdets, which we analyze separately to isolate their impact: first with adaptive data generation but no shared limit, then with a shared limit but no data adaptivity. In both cases we assume no intentional collusion, yet show by counterexample that per-querier privacy loss grows linearly with the number of queriers. We then identify a strong siloing assumption under which per-querier enforcement recovers soundness.
3.1  Formal model of multi-querier system
Alg. 1 defines our generic model of a system with $K$ queriers ($Q^{1},\dots,Q^{K}$), each with its own privacy budget $\mathcal{F}^{k}$ and interacting with a shared database $D$ over discrete time steps. The model accommodates both streams and static databases: a data generation process $\mathcal{G}$ may produce new data at each step (appended to $D$), while in the static case $\mathcal{G}$ outputs nothing after the initial step. It also supports both per-querier enforcement, via $\mathcal{F}^{k}$, and an abstract shared limit $\mathcal{L}$, which can represent a global budget. At each step, querier $k$ selects a query $Q^{k}_{t}$ based only on its own past results; this excludes intentional collusion by construction. The query is executed if both $\mathcal{F}^{k}$ and $\mathcal{L}$ approve; if so, the querier receives the new result $V_{t}^{k}$, which is appended to its own view of the database. Otherwise, the querier receives an empty result $\bot$ that is independent of the database.
To analyze DP for such a system, one introduces a challenge bit $b$ that governs whether a particular record $x_{0}$ is added to the database at time $t_{0}$. Neither queriers nor $\mathcal{G}$ see $b$, but as queriers run more queries, their views may reveal information about it – thus modeling the increased “leakage” of an individual’s presence through the results of queries on a DP database. Privacy loss quantifies how hard it is for an adversary to distinguish between $b=0$ and $b=1$, whether the adversary controls a single querier (per-querier guarantees) or all queriers (global guarantees).
Formally, Alg.  1 satisfies $\epsilon$-DP for querier $k\in[K]$ if
$\forall x_{0}\in\mathcal{X},\forall t_{0}\in[t_{\max}],\forall\mathcal{G},\forall(\mathcal{Q}^{1},\dots,\mathcal{Q}^{n+1}),$
$\forall S\subseteq\mathcal{R}^{t_{\max}},\left|\ln\left(\frac{\Pr[V^{k,(0)}\in S]}{\Pr[V^{k,(1)}\in S]}\right)\right|\leq\epsilon.$
By contrast, a global guarantee would bound $\frac{\Pr[V^{(0)}\in S]}{\Pr[V^{(1)}\in S]}$, if for example $\mathcal{L}$ were implemented as a global budget.
We next analyze per-querier guarantees and show two orthogonal cases where they degrade linearly with the number of queriers. These correspond to the red conditions in Alg.  1: the first models adaptive data generation in streaming settings, the second models the presence of a shared limit in addition to per-querier budgets. By toggling these conditions on/off, we study the impact of each condition on per-querier guarantees in isolation. Appendix A.3 extends these results to the context of individual DP budgets.
Alg. 1 Formal model of generic multi-querier, pure-DP system with per-querier budgets.
1:Input
2: Challenge bit $b\in\{0,1\}$
3: In-or-out record $x_{0}$ to be inserted at time $t_{0}$
4: Data generation process $\mathcal{G}$
5: Queriers $\mathcal{Q}^{1},\dots,\mathcal{Q}^{K}$
6: Per-querier budgets, implemented as pure-DP filters $\mathcal{F}^{1},\dots,\mathcal{F}^{K}$
7: Shared limit $\mathcal{L}$
8:for $t\in[t_{\max}]$ do
9: // Data generation model:
[…]
11:if $t=t_{0}$ and $b=1$ then // insert in-our-out record on challenge
[…]
13: // Query execution model:
14:for $k\in[K]$ do
[…]
16:if $\mathcal{F}^{k}(q_{1}^{k},\cdots,q_{t}^{k}){\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\wedge\mathcal{L}(q_{1}^{1},\cdots,q_{1}^{K},\cdots q_{t}^{k})}$ then // query is executed only if both $\mathcal{Q}^{k}$’s budget and shared limits permit it.
17:  $V^{k}_{t}\leftarrow q_{t}^{k}(D_{\leq t})$
18:else
19:  $V^{k}_{t}\leftarrow\bot$
20:return $V^{1},\dots,V^{K}$
3.2  Impact of adaptive data generation
We first consider the case where data is generated over time with no shared limits ($\mathcal{L}$ always outputs TRUE). Adaptivity in data generation – where new data depends on past DP query results – has been studied in the context of streaming algorithms (GS13, 18, 9, 21) and interactive DP systems (sage, 25, 29, 41, 16). Modeling data adaptivity is essential: (1) it directly impacts DP analyses, as these works and ours show, and (2) it is inevitable in many streaming systems. For instance, a movie recommender may use prior DP measurements to suggest a film, and the user’s decision to watch it becomes new data for future measurements. In advertising, such feedback loops are also pervasive.
Line 9 of Alg.  1 captures a general form of data adaptivity for multi-querier systems: $D_{t}\leftarrow\mathcal{G}(V_{<t}^{1},\dots,V_{<t}^{K})$, meaning data can depend not only on one querier’s past results but also on others’. This cross-querier dependency arises even without collusion, through user behavior or mechanisms like auctions. For example, news.ex may show an impression for either shoes.ex or hats.ex, depending on bids shaped by prior query results from both sites.
We prove that under this model, per-querier budgets (formalized as privacy filters (RRUV16, 37)) with capacity $\epsilon$ do not suffice:
Theorem 1 (Per-querier guarantees are unsound under adaptive data generation).
Consider Alg. 1, with $K=n+1$ queriers, where each $\mathcal{F}^{k}$ is a pure-DP filter with capacity $\epsilon>0$, and where $\mathcal{L}$ always outputs TRUE (i.e., no shared limit). Denote by $V^{n+1,(b)}$ the view of $\mathcal{Q}^{n+1}$ on challenge bit $b$. Then, there exists $\epsilon^{\prime}>0$ such that:
$\exists x_{0}\in\mathcal{X},\exists t_{0}\in[t_{\max}],\exists\mathcal{G},\exists\mathcal{Q}^{1},\dots,\mathcal{Q}^{n+1},$
$\exists S\subseteq\mathcal{R}^{t_{\max}}:\left|\ln\left(\frac{\Pr[V^{n+1,(0)}\in S]}{\Pr[V^{n+1,(1)}\in S]}\right)\right|\geq\epsilon+n\epsilon^{\prime}.$
Thm. 1 shows that even with per-querier budgets of capacity $\epsilon$, privacy loss can exceed $\epsilon$, with excess loss growing linearly in the number of queriers. This arises not from collusion but from worst-case analysis that DP must account for. The proof, given in Appendix A.1, works by constructing an example of data generation and querying processes that exhibits that much excess loss. While the theorem addresses only pure-DP guarantees (yielding a linear increase in excess loss), the result is likely to extend to approximate DP, where excess loss would grow with the square root of the number of queriers (proof is in progress).
Prior analyses overlooked this point. ARA’s implementation supports multiple adtechs (Goo22, 17), but its proof models only a single querier (GHH+25, 16), effectively defining data generation as $\mathcal{G}(V^{k}_{<t})$ for a fixed $k$ rather than our more general $\mathcal{G}(V_{<t}^{1},\dots,V_{<t}^{K})$. Cookie Monster, by contrast, allowed multiple queriers but modeled a static dataset in its proof (TKM+24, 41).
3.3  Impact of shared limits
We next consider the case where a shared limit is added to a system with per-querier enforcement (line 16 in Alg. 1). Such limits have been proposed in both static (GHK+18, 14, 43) and streaming (KVLH25, 24) DP systems as defense-in-depth to bound privacy loss under collusion. We instantiate $\mathcal{L}$ as a pure-DP filter shared across all queriers, which lets us quantify the excess privacy loss. The same reasoning, however, applies to other shared-limit mechanisms, such as a cap on the total number of queries, or rate limiters or quotas spanning querier subsets.
Using a proof technique similar to the adaptive data case, we show that even without data adaptivity, per-querier privacy loss can exceed $\epsilon$ when a shared budget is present. To separate this result from Thm. 1, we focus on static databases (restricting $\mathcal{G}$ to return empty set after the first time step). As before, the excess loss grows linearly with the number of queriers sharing the limit (proof in Appendix A.2).
Theorem 2 (Per-querier guarantees are unsound under adaptive query budgets with shared limits).
Take $n+1$ queriers, with per-querier filters with capacity $\epsilon$, and a shared filter with capacity $\epsilon_{g}<(n+1)\epsilon$. There exists $\epsilon^{\prime}>0$ such that for all $n\geq\frac{4\ln 2}{(1-e^{-\epsilon/2})^{2}}$, there exists $\mathcal{G}$ that returns $\emptyset$ after the first time step, such that
$\exists x_{0}\in\mathcal{X},\exists t_{0}\in[t_{\max}],\exists\mathcal{Q}^{1},\dots,\mathcal{Q}^{n+1},$
$\exists S\subseteq\mathcal{R}^{t_{\max}}:\left|\ln\left(\frac{\Pr[V^{n+1,(0)}\in S]}{\Pr[V^{n+1,(1)}\in S]}\right)\right|\geq\epsilon+n\epsilon^{\prime}.$
This effect has been missed in prior work on multi-granularity enforcement, either because it was not formalized at all (GHK+18, 14) or because per-querier and global limits were analyzed in isolation or under non-adaptive queriers (KVLH25, 24, 43). Conceptually, this is like treating the two conditions in $\mathcal{F}^{k}(q_{1}^{k},\cdots,q_{t}^{k})\wedge\mathcal{L}(q_{1}^{1},\dots,q_{1}^{K},\dots,q_{t}^{k})$ as if they acted separately, when in real systems they inevitably apply together.
3.4  Siloing assumption
It is interesting to ask whether there exist conditions – realistic or not – under which per-querier enforcement can be made sound. We identify one such condition, which we call the siloing assumption, and prove in Appendix A.4 that it suffices to recover per-querier DP guarantees in our general model with both data adaptivity and shared limits (Alg. 1). While this assumption is strong – likely too strong for any real system – it is useful for clarifying the essence of per-querier semantic limitations.
Theorem 3 (Siloing assumption).
Take $K$ queriers, $k\in[K]$, a per-querier filter $\mathcal{F}^{k}$, with capacity $\epsilon>0$, and a shared limit $\mathcal{L}$. Suppose that the two following assumptions hold:
  1. (1)
    $D_{t}\leftarrow\mathcal{G}(V^{1}_{<t},\dots V^{K}_{<t})$ is of the following form: $D_{t}^{1}\leftarrow\mathcal{G}(V^{1}_{<t}),\dots,D_{t}^{K}\leftarrow\mathcal{G}(V^{K}_{<t})$, and $q_{t}^{k}(D_{\leq t})=q_{t}^{k}(D_{\leq t}^{k})$.
  2. (2)
    Shared limits do not reject $k$’s reports depending on other queriers’ results:
    $\forall k\in[K],\forall u,w\in\mathcal{R}^{t}:u^{k}=w^{k},$
    $\mathcal{L}\left(q_{1}^{1},\dots,q_{1}^{K},\dots,q_{t}^{1}(u^{1}_{<t}),\dots q_{t}^{k}(u^{k}_{<t})\right)$
    $=\mathcal{L}\left(q_{1}^{1},\dots,q_{1}^{K},\dots,q_{t}^{1}(u^{1}_{<t}),\dots q_{t}^{k}(u^{k}_{<t})\right)$
Then,
$\forall x_{0}\in\mathcal{X},\forall t_{0}\in[t_{\max}],\forall\mathcal{G},\forall\mathcal{Q}^{1},\dots,\mathcal{Q}^{K},$
$\forall S\subseteq\mathcal{R}^{t_{\max}},\left|\ln\left(\frac{\Pr[V^{k,(0)}\in S]}{\Pr[V^{k,(1)}\in S]}\right)\right|\leq\epsilon.$
Intuitively, the siloing assumption requires that (1) each querier operates on a disjoint data stream, and (2) shared limits either never bind or are configured so loosely that they never fail on any device. These are heavy assumptions, unlikely to hold in practice. We strongly recommend not trying to engender them – for example, by disabling or weakening shared limits – as doing so would undermine the very goal of providing meaningful user privacy guarantees.
Instead, our message is three-fold. First, Attribution should explicitly acknowledge the semantic limitations of per-querier enforcement; see our proposed specification text (bens-pr, 2). Second, we recommend augmenting per-querier budgets with global enforcement, and our second contribution – and remainder of this paper – shows how to do this robustly and efficiently. Third, we urge the DP community to raise the bar on system modeling fidelity – a point long emphasized in formal methods, but not yet as ingrained in DP as it should be. We have seen the pitfalls repeatedly: prior ARA-related analyses mis-modeled its budgeting, Cookie Monster and later ARA work overlooked cross-querier data adaptivity, and other DP papers ignored multi-granularity interactions. Per-querier enforcement is just one case study in a broader lesson: modeling fidelity is as critical as the mathematics of DP itself.
4  Contribution 2: Global IDP with Big Bird
The preceding section revealed semantic gaps in per-querier enforcement, arising from both intentional collusion and inherent effects such as cross-querier data adaptation and shared limits. Still, per-querier budgets remain an important part of a DP system like Attribution: because each querier’s budget is isolated, they can be configured tightly (with small $\epsilon_{\textnormal{querier}}$) without risk of exhaustion by others. By contrast, global enforcement – such as through a shared budget across all queriers – offers sound semantics, holding under both collusion and the effects that invalidate per-querier guarantees. The tradeoff is twofold: (1) the global budget must be set more loosely to sustain a workload of multiple queriers (large $\epsilon_{\textnormal{global}}$), and (2) it introduces budget depletion risks since many queriers compete for the same shared budget.
For this reason, the Attribution specification includes the notion of safety limits spanning all queriers but leaves their enforcement undefined (ppa-spec-safety-limits, 39), with an open concern about denial-of-service (DoS) attacks that deplete globally shared budgets (charlie-issue, 38). In this section, we present Big Bird, the first design for enforcing a global privacy guarantee while incorporating resilience against DoS budget depletion attacks. While Big Bird is developed specifically for Attribution, where it enforces device-epoch-level IDP globally across all queriers, it also provides a blueprint for other DP systems that rely on global budget enforcement but may deviate from Attribution’s specific privacy unit or privacy definition.
Enforcing the global privacy guarantee is central to Big Bird’s mission, and we provide proof of its soundness in the context of a formal system model that exhibits a much higher degree of fidelity than is typical in DP analyses (§4.8). However, the bulk of this section focuses on depletion attacks and the resilience we incorporate in Big Bird against them. We start this focus by defining the threat model for these attacks.
4.1  DoS depletion threat model
(a) Sybils as conversion sites
(b) Sybils as intermediaries
Fig. 2. Depletion attack examples. Sybil queriers deplete the global budget from a single user action, by acting as (a) conversion sites or (b) intermediaries. Per-querier and global budgets are depicted as batteries (filled = available, empty = drained).
Global budgets are a shared resource, so under Attribution’s open-web design, any site – benign or malicious – can invoke the API and draw from them. An attacker can exploit this by registering many Sybil domains, which are cheap on the web, and coordinating them to deplete a user’s global budget. If the attacker succeeds across many devices, legitimate sites will be locked out of measurement: a distributed DoS at the privacy layer. Motivations vary (undermining the API indiscriminately, blocking particular queriers, or improving the attacker’s own utility without regard for others), but we focus on the most damaging case – indiscriminate DoS – because it is both feasible under the current Attribution spec (which defines per-querier and global budgets but provides no defenses against depletion) and especially destructive to the API’s utility.
Fig. 2(a) shows an aggressive instance. When a user visits malicious site x.ex, the site registers impressions under many Sybil domains that pose as conversion sites and then auto-redirects the user through them. Each Sybil calls measureConversion(), consuming global budget like any legitimate querier. Because each request passes through per-querier budget checks, each Sybil can spend up to $\epsilon_{\text{querier}}$. With roughly $\epsilon_{\text{global}}/\epsilon_{\text{querier}}$ Sybils, x.ex can exhaust that user’s global budget from a single visit. Consequently, when the user later visits a real merchant (e.g., y.ex), that merchant receives null reports – both useless and biased. If x.ex can attract many users to visit its site even once each, it can disable Attribution for honest queriers across the epoch (typically a week), creating long-term unavailability of the API.
Attacks need not be so overt. Instead of redirections, an attacker could embed many Sybil intermediaries (Fig. 2(b)), each requesting reports and drawing from distinct per-querier budgets – a vector that becomes relevant once Attribution supports ad-techs as queriers. Even subtler, an attacker can pace Sybil activity across visits and users to slowly drain budgets, making malicious behavior harder to distinguish from benign load.
Clearly, embedding resilience to such attacks is critical for Attribution’s success. The key question is how to achieve this without undermining the API’s openness, flexibility, or utility. In §4.2, we explore several natural – but flawed – approaches that highlight the tension between protecting the global budget and preserving flexibility for benign use. These motivate the design principles underlying our own approach in Big Bird (§4.3). Before turning to those, we summarize the assumptions and goals that shape our design.
Assumptions. Our strategy is to embed resilience directly into the global budget enforcement layer, with the explicit goal of constraining malicious workloads to behave within the bounds of expected benign use. Rather than trying to detect or block Sybils, Big Bird makes them unscalable: attackers can still operate but must pay a cost proportional to genuine user engagement. We assume domain creation is effectively free for attackers, and measure attack cost instead in terms of the number of intentional user actions (e.g., visits or clicks) the attacker must induce across enough devices to significantly degrade measurement accuracy for a non-trivial fraction of benign queriers. Critically, we assume that adversaries cannot easily induce large numbers of users to intentionally visit many distinct Sybil domains, and that browsers can distinguish intentional user actions from automated navigations. If these assumptions fail, Big Bird still preserves global IDP guarantees – privacy never breaks – but its resilience to DoS degrades.
We exclude from scope attacks that involve malicious bots submitting fabricated reports (including null reports) to poison aggregate measurements for legitimate queriers. Such poisoning attempts target data quality rather than global-budget depletion. While these must eventually be mitigated within Attribution, our focus here is on preserving availability of measurement of benign-user data by legitimate queriers – a prerequisite for meaningful defenses against poisoning as well, since maintaining sufficient volume of benign-user data is essential to dilute the effect of adversarial noise.
Finally, we assume that benign sites – the ones Big Bird seeks to protect – do not partner, directly or transitively through their partners, with malicious sites, for example by embedding them as intermediaries or serving their ads. This reflects a limitation of our vantage point: Big Bird enforces resilience at the level of global budget management, not partner vetting or ad-network governance. Coordinated attacks that exploit partnership chains require complementary ecosystem defenses such as registration, vetting, or auditing. Big Bird instead focuses on protecting benign sites unconnected to adversaries but at risk of losing API access due to global-budget depletion by them. In short, adversarial impact should remain localized to sites that directly or transitively interact with Sybils, while harm to unrelated benign sites should remain bounded.
4.2  Candidate approaches
One can imagine many restrictions or limits that could be added to Attribution to guard the global budget against depletion attacks. Each such restriction faces the same dilemma: effectiveness against attacks must be balanced against the risk of over-constraining benign use. For an API still in its infancy, premature restrictions can be especially harmful: limits placed on the wrong quantity may stifle future use cases. Worse, they can leave capacity unused – budget left on the table – when a misaligned limit triggers before the global budget is truly exhausted.
Rate-limit API calls. One could cap the number of impressions a site may log or the number of reports it may request. But these are indirect metrics: API calls do not themselves consume privacy budget. Some sites make many calls yet consume little, while others consume heavily with only a few. This is particularly true under IDP, where consumption depends on a user’s actual data contribution. For example, most impressions never lead to conversions, and many conversions occur without any attributable impression (e.g., people often buy shoes without seeing an ad, simply because they need them). Such events never consume global budget under Attribution’s IDP accounting, yet would still count against API-call caps. Hard limits on such misaligned metrics risk penalizing benign workloads, wasting global budget, or leaving attackers free to operate below thresholds that are loosely configured precisely due to this misalignment.
Rate-limit global budget consumption. A more direct approach is to constrain budget consumption itself, for example by limiting how much of the global budget can be spent within short timeframes. This would disable the aggressive attacks in Fig. 2, which rapidly burn budget, but fails against gradual strategies that spread consumption over time. For instance, if a user merely leaves the x.ex tab open, the site could pace its global-budget use via auto-redirects or intermediary requests to bypass short-window rate limits. Extending the rate-limit window to capture such attacks risks harming benign workloads or make the limiter itself a DoS target.
Cap the count of per-querier budgets. Since Attribution already assigns per-querier budgets, one could cap how many such budgets can exist on a device. For instance, only the first $N$ domains to request reports based on an epoch’s data would receive budgets for that epoch; later domains would always see null reports. If NN $\geq\epsilon_{\textnormal{global}}/\epsilon_{\textnormal{querier}}$, the cap is ineffective; if more restrictive, it simply becomes a new DoS target.
Cap global budget consumption per user action. The striking feature of Fig. 2 is that a single user action lets x.ex drain the user’s entire global budget. A natural fix is to cap budget use per user action: when a user visits x.ex, a budget of $\epsilon_{\text{action}}$ is initialized; any queriers in the ensuing redirect chain or embedded in the site can together consume up to this cap, after which reports are blocked until the next user action. This ties budget use more closely to user actions – the costly resource in our threat model – yet still allows a single high-engagement site to deplete budget over many interactions. Still, among candidate fixes, this approach most closely anticipates the direction we take in Big Bird.
4.3  Principles and stock-and-flow insight
(a) Privacy loss in normal use
(b) Privacy loss in attack
Fig. 3. Stock-and-flow pattern of global-budget consumption in Attribution. (a) Normal use is expected to follow this pattern, where user actions on impression sites creates stock and user actions on conversion sites trigger flows. (b) Attacks break this pattern by automatically creating many stocks and triggering flows from very few user actions.
The preceding approaches highlight the core challenge: how to defend the global budget without crippling legitimate use. Our design of Big Bird is guided by three principles. P1: Where possible, limits should align with actual budget consumption, not indirect proxies like API call counts, to avoid wasting capacity. P2: Defenses should target structural properties of benign use, rather than specific attack tactics (such as rapid burn or the use of redirects or embedded intermediaries), so they remain effective with new, unforeseen attack shapes. P3: Budget consumption should be ultimately tied to intentional user actions across sites – since that is what the Attribution API is meant to measure in the first place. These principles lead directly to our key insight behind Big Bird: in normal operation, privacy loss in Attribution can be expected to follow a predictable stock-and-flow pattern, which can be enforced to provide resilience against depletion attacks.
Stock-and-flow. Fig. 3 illustrates this pattern. Specifically, the Attribution API is intended to measure user interactions across distinct sites, such as impression and conversion sites. In normal Attribution use (Fig. 3(a)), privacy loss should only be consumed when there is something to measure: a user visits one site (and sees an ad) and later visits a distinct site (and “converts” there). That creates a stock-and-flow pattern for privacy budget consumption: (1) stock is created when the user visits an impression site and the site saves an impression; and (2) flow is triggered when a user visits a conversion site and the site requests a report. Non-zero privacy loss is realized only when both steps occur, a property afforded by the IDP mechanism. By contrast, attacks disrupt this structure (Fig. 3(b)) by coalescing domains or automating stock creation and flow triggering – through automatic redirections, embedded intermediaries, or similar techniques – allowing substantial global budget consumption from minimal user actions. Big Bird restores this pattern by putting privacy-loss-based quotas both on stock creation and on flow triggering, and ultimately tying these to intentional user actions.
4.4  Big Bird architecture
Fig. 4. Big Bird architecture. Enforces stock-and-flow quotas at impression and conversion sites, and limits the creation of new quota budgets per user action. Depicts quota budget structure for a single epoch, though in reality quota budgets exist for each epoch.
Fig. 4 shows the Big Bird architecture. It sits as a management layer around Attribution’s global budgets and protecting them from depletion by enforcing quotas that restore the expected stock-and-flow pattern: (1) impression-site quotas cap how much stock any one domain can contribute; (2) conversion-site quotas cap how much flow a domain can trigger; and (3) a quota-count cap limits how many distinct new quota budgets of each type can be created in the context of a single user action.
The first two quotas are privacy budgets implemented as IDP filters, just like global and per-querier budgets. However, their purpose is not privacy; they protect the global budget from depletion. These quotas account for individual privacy loss in a manner aligned with the global budget (principle P1), and constrain stock creation and flow triggering according to the expected structure of benign behavior (principle P2). By contrast, the quota-count cap is not a privacy budget; it simply restricts the number of new quota budgets created, under different domain names, from a single user action. Together, these three quotas constrain consumption of the global budget, forcing the attacker to behave closer to the expected workload, and tie this restricted consumption to something tangible, such as intentional user actions across sites (principle P3).
To illustrate, consider the example workflow in Fig. 4: 1 a user sees a shoe ad on news.ex and 2 another on blog.ex; they purchase on shoes.ex, which 3 requests two reports from Big Bird– one, labeled (i), for its own measurement to compare the efficacy of the two shoe-ad creatives, and the other, labeled (j), for placement optimization by its intermediary, adtech.ex. 4 Big Bird generates these reports much like the Attribution API: it selects matching impressions, checks and accounts for privacy loss, and generates the capped histogram, but it uses a different privacy accounting algorithm (step (4.2)). Detailed in the next section, this algorithm deducts – atomically across all budgets – a per-querier budgets (shoes.ex, adtech.ex); b the global budget; c impression-site quotas (news.ex, blog.ex); and d the conversion-site quota (shoes.ex). It also e enforces the quota-count cap, ensuring no excessive number of new quotas are instantiated following a single user action. In this example, each user action creates only one new quota – one from visiting news.ex, one from blog.ex, and one from shoes.ex – so the reports proceed. If any check fails, null reports are returned and no budgets are deducted; otherwise, all deductions are applied atomically.
Finally, 5 two encrypted reports are returned to the ad-tech: one (i) to be used to conduct shoes.ex’s measurement of the two shoe-ad creatives and one (j) to be used for the ad-tech’s own placement optimization. For this example, these two reports contain the same histogram attributing the conversion value to each of the two impressions the user has seen, but they are deducted separately from the global budget and from the impression-site and conversion-site quota budgets, as two separate queries will be run on each. However, we discover situations in which different reports released for a single conversion need not deduct separately from the global budget or the quotas – specifically if they contain different and non-overlapping portions of the histogram. This is an opportunity for privacy budget optimization that we prove for the first time in this paper and is important for efficient management of the global budget.
While the architecture may appear “simple” at a high level – a plus for deployability – several challenges arise when instantiating it, which we address in the remainder of this section: (§4.5) enforcing the quotas correctly and efficiently, without double-counting privacy loss against shared budgets; (§4.6) configuring quotas to support benign workloads, ensure efficient global budget utilization, while at the same time constraining attacker behavior; (§4.7) assessing the resilience properties these quotas afford against depletion attacks ; and (§4.8) proving global device-epoch-level IDP for the end-to-end design. All of these constitute substantial and difficult contributions, both for the Attribution community and for the systems and DP research communities.
4.5  Big Bird algorithm
Big Bird extends the Attribution API with new enforcement logic for its three quota types—impression-site, conversion-site, and quota-count—while maintaining device–epoch–level IDP. It introduces one new API function, onUserAction() and augments two existing ones, saveImpression() and measureConversion(). Alg. 2 illustrates all three. Together, these functions link global-budget consumption to intentional user actions and enforce the stock-and-flow structure that yields Big Bird ’s resilience guarantees.
User actions as the root of quota creation. Each time a genuine user interaction is detected, the browser calls the new onUserAction() function, which initializes a new user action context. Within this context, Big Bird tracks which top-level domains invoke the API through an accessedSites set that resets on every new user action, ensuring that subsequent quota creation and budget consumption are explicitly linked to user-driven behavior (principle P3).
Quota-count cap. Within each user-action context, Big Bird enforces a quota-count cap that limits how many distinct impression or conversion sites – identified by their top-level eTLD+1 domains – can create new quota budgets in response to that action. Unlike naïve rate limits on API calls, this cap does not restrict invocation frequency; it bounds the number of distinct quota budgets that can be instantiated per intentional user action. This distinction is central to principle P1: whereas API-call limits regulate an indirect proxy for privacy loss, the quota-count cap constrains the creation of budgets that will account for actual privacy loss when reports are later generated. Crucially, it still allows an unbounded number of impressions to be registered or reports to be requested for a given user action – as long as these occur within a small set of top-level domains. This structurally blocks amplification attacks such as redirection chains, where Sybil domains rotate to drain budgets, without impeding benign workloads, which typically involve only one or a few domains saving impressions or requesting reports per user action. For example, a user clicking a link in an email or shortened URL may trigger a legitimate redirection, but under normal operation the destination site will not continue redirecting while invoking the Attribution API from new domains. A small quota-count value (e.g., 1–2) enforces this behavioral boundary while preserving compatibility with legitimate redirections, which are too fundamental to the web to prohibit outright.
Impression-site quotas. When an impression is logged via saveImpression(), the corresponding site is added to the accessedSites set. If adding the site would exceed the quota-count cap, the impression is rejected. Otherwise, the impression is recorded in the current epoch’s database. Each impression site maintains its own imp-quota– a privacy-loss–based quota that tracks how much global-budget stock it can contribute in that epoch. This quota operates at the device-epoch-impression-site level (finer than the device–epoch granularity of other filters). This is because while multiple impression sites can be requested for a single conversion request, most requested sites might not have displayed any relevant impressions for that conversion. Accounting at the device-epoch-impression-site level allows such impression sites to not deduct any budget.
Conversion-site quotas. When a conversion occurs, measureConversion() adds the conversion site to accessedSites, again subject to the quota-count cap. Big Bird then identifies all matching impressions within the attribution window, computes per-epoch individual privacy losses, and then, on an epoch basis, it atomically verifies that all relevant budgets – querier, global, impression-site, and conversion-site – can jointly accommodate those losses. If any budget lacks capacity, the epoch’s data is dropped from consideration and no deductions occur, preventing partial or wasted consumption. This is done by wrapping all privacy budgets’ and quotas’ consumptions in each epoch into an atomic transaction with serializable isolation.
Stock-and-flow pattern. The quotas we introduce in Attribution ensure that privacy loss is constrained to follow the stock-and-flow pattern – a structural property expected for benign use of the API (principle P2). The impression-site budget constrains stock by ensuring no single impression site can contribute too much, and the conversion-site budget constrains the flow occuring through any given conversion site. These two, along with the quota-count cap, are the roots of the resilience guarantees we prove for Big Bird in §4.7.
Privacy loss accounting example. We illustrate the Big Bird algorithm with the example from Fig. 4 (which is similar to the one from Fig. 1), where a user saves 1 an impression from the first creative on news.ex at epoch e1, and 2 an impression from the second creative on blog.ex at epoch e2. The user then 3 makes a purchase on shoes.ex, which 4 calls measureConversion twice, with parameters of the form impressionSites=[blog.ex, news.ex], intermediarySites=[adtech.ex], epsilon=0.5, value=60, maxValue=100, histogramSize=3), first with shoes.ex as a querier (report (i) for measurement purposes) and then with adtech.ex as a querier (report (j) for placement optimization). Finally, 4 generates an attribution histogram for each querier, after accounting for privacy loss for both the reports.
Privacy loss accounting for (i), for which shoes.ex is the querier, happens as follows. First, Big Bird computes the device-epoch privacy loss for each epoch. Since e1 and e2 both contain a relevant impression, by Cookie Monster’s algorithm their privacy loss is $\mathsf{(value/maxValue)}\cdot\mathsf{epsilon}=0.3$. Meanwhile, e3 has no privacy loss. Big Bird also computes the device-epoch-impression-site privacy loss: (e1, news.ex) and (e2, blog.ex) both have 0.3 privacy loss, while other epoch-impression site pairs on the device have no privacy loss. Second, Big Bird checks whether each budget can deduct the corresponding privacy loss, for each epoch. For e1, the global budget is able to deduct 0.3. In that example, shoes.ex is simultaneously conversion site and querier, and both budgets are able to deduct 0.3. The impression site budget for (e1, news.ex) is also able to deduct 0.3. Accounting is similar for e2, and e3 has no privacy loss. Third, since all the checks pass, the transaction is executed atomically across budgets within each epoch, and impressions from e1 and e2 are used to generate the attribution histogram.
Accounting works similarly for (j), where the ad-tech is the querier – this time deducting from adtech.ex’s querier budget. Table 1 shows the state of the budgets after executing (i) and (j) (adtech2.ex is introduced below).
epoch querier $\epsilon_{\textnormal{querier}}=1$ global $\epsilon_{\textnormal{global}}=8$ imp-quota $\epsilon_{\textnormal{imp-quota}}=4$ conv-quota $\epsilon_{\textnormal{conv-quota}}=2$
shoes.ex adtech.ex adtech2.ex news.ex blog.ex shoes.ex
e1 0.7 0.7 0.7 7.4 3.4 4 1.4
e2 0.7 0.7 0.7 7.4 4 3.4 1.4
e3 1 1 1 8 4 4 2
Tab. 1. Privacy Budget Store in Big Bird. Remaining capacities of all privacy and quota budgets after executing the example in Fig. 4, as described in the privacy-loss accounting example of §4.5. All budgets are assumed to be fully charged to their initial capacities (shown in the header) prior to execution.
Avoiding double-counting privacy loss against shared budgets. It is common for conversion sites to embed multiple intermediaries, and to perform attribution once across impressions displayed by all intermediaries. When multiple reports for the same conversion involve disjoint impression sets (e.g., across intermediaries), we prove that one need not double-count against shared budgets (global, conversion site quota, impression site quota). Indeed, the combination of reports from different intermediaries reveals no more than a single attribution across intermediaries. We present the intuition with an example here; detailed formalism and proofs are in Appendix D.
We modify the example as follows. Instead of working with a single intermediary adtech.ex, shoes.ex works with both adtech.ex and adtech2.ex. adtech.ex displays impressions on news.ex, and adtech2.ex on blog.ex. Upon conversion, attribution is still computed across all relevant impressions, spanning both intermediaries. For shoes.ex’s own measurement queries, accounting is unchanged since intermediaries operate on shoes.ex’s behalf and are not queriers themselves. As before, adtech.ex can request an additional report (j) measuring its own impressions for placement optimization purposes. Here, adtech.ex receives a report containing a histogram with attributed value 30 (for the impression on news.ex). Additionally, adtech2.ex can measure its own impressions and receive a report (k) containing a histogram with attributed value 30. Naively, one would expect a cumulative deduction of 0.9 from the global budget (three reports). Yet, one can observe that the two intermediary reports are shards of a single report with attributed value 60, because adtech.ex and adtech2.ex receive disjoint bins of the histogram. Leveraging this observation, we show that it is possible to deduct only 0.6 from the global budget. The same optimization applies to shared quotas.
Alg. 2 Big Bird Algorithm
def onUserAction(uaCtx)
  // Set of top-level sites accessed in the context of the
  // current user action (uaCtx).
  accessedSites[uaCtx] $\leftarrow\emptyset$
def saveImpression(impSite, impression, uaCtx)
  if $|\text{accessedSites[uaCtx]}\cup\text{\{impSite\}}|>k_{\textnormal{quota-count}}$ then
   return // can’t allow new imp-quotas from this user action
  $\text{accessedSites[uaCtx]}\ \cup=\text{\{impSite\}}$
  ImpressionStore[current epoch].add(impression)
  return
def measureConversion(impSites, querier, convSite,
   reportRequest, uaCtx)
  if $|\text{accessedSites[uaCtx]}\cup\text{convSite}|>k_{\textnormal{quota-count}}$ then
   // Can’t allow new conv-quotas from this user action.
   return AttributionFn(,,)($\emptyset$$\dots,\emptyset$) // return null report
  $\text{accessedSites[uaCtx]}\ \cup=\text{\{convSite\}}$
  // Separately for each epoch, match impressions and do privacy
  // loss accounting (Steps (4.1) and (4.2) in Fig. 4).
  for $e\in\{e_{1},\dots,e_{w}\}$ do // epochs in attribution window
   $I_{e}\leftarrow$ MatchingFn(ImpressionStore[e], reportRequest)
   $\epsilon_{e}\leftarrow$ computeEpochLoss($I_{e}$, reportRequest)
   for $i\in$ impSites do
  $\epsilon_{e}^{i}\leftarrow$ computeEpochImpSiteLoss($I_{e}[i]$, reportRequest)
   // Execute budget checks & deductions as a transaction.
   tx = beginTransaction()
   if querierBudget[e][querier].tryConsume($\epsilon_{e}$, tx) and
   globalBudget[e].tryConsume($\epsilon_{e}$, tx) and
   convQuota[e][convSite].tryConsume($\epsilon_{e}$, tx) and
   (impQuota[e][$i$].tryConsume($\epsilon_{e}^{i}$, tx) $\forall i\in$ impSites) then
  tx.commit() // all budgets deducted successfully
   else // a budget deduction failed
  tx.abort() // undo any other budget deductions
  $I_{e}\leftarrow\emptyset$ // drop epoch’s impression data from consideration
  // Generate attribution histogram from remaining impressions
  // (Step (4.3) in Fig. 4).
[…]
4.6  Quota configuration
How should filter and quota capacities be configured? There are two challenges. First, each budget should be configured as tightly as possible to prevent attacks without disrupting benign workloads. Second, the budgets as a whole should be consistent to avoid tripping each other up or being unnecessarily loose. We approach these challenges from a utility-first perspective, by defining and solving a constrained optimization problem as follows.
First, we take the querier budget $\epsilon_{\textnormal{querier}}$ as a configuration parameter, which represents what is needed to support a single querier’s expected workload. Next, we express all the other budgets as functions of $\epsilon_{\textnormal{querier}}$, thanks to browser-adjustable parameters describing expected workload scale:
  • On any conversion site, the conversion site itself should be able to use its full querier budget for its own queries. Additionally, intermediaries should be able to issue their own queries, up to some fraction of the conversion site’s own queries. Hence, we take a variable $r$ representing the expected share of budget for intermediaries, and have the following constraint: $\epsilon_{\textnormal{conv-quota}}\geq(1+r)\epsilon_{\textnormal{querier}}$ (to support $\epsilon_{\textnormal{querier}}$ for single-advertiser queries plus $r\epsilon_{\textnormal{querier}}$ for cross-advertiser queries).
  • The global budget should not be completely depleted by flow through a single conversion site. We define $N$, the number of conversion sites that need to be supported from a single epoch, which gives $\epsilon_{\textnormal{global}}\geq N\epsilon_{\textnormal{conv-quota}}$.
  • Similarly, the global budget should be larger than a single impression site’s stock. We take an expected number of impression sites $M$ and get $\epsilon_{\textnormal{global}}\geq M\epsilon_{\textnormal{imp-quota}}$.
  • Finally, multiple conversion sites request impressions from the same impression site, so a small $\epsilon_{\textnormal{imp-quota}}$ could prevent conversion sites from using their full $\epsilon_{\textnormal{conv-quota}}$. A parameter $n$ allows an impression site to be queried by $n$ conversion sites, with $\epsilon_{\textnormal{imp-quota}}\geq n\epsilon_{\textnormal{conv-quota}}$.
Finally, we solve the problem to minimize $\epsilon_{\textnormal{global}}$, resulting in capacities presented in Table 2. This ensures Big Bird provides the strongest privacy guarantees possible while satisfying our utility constraints.
The parameters can be configured from expected workload characteristics, which can be derived periodically from simple, robust statistics. Our evaluation (§6.3) provides an example, such as the 95th percentile of the number of conversion sites requesting an epoch, across devices. Our experiments show that Big Bird’s performance is stable across a wide range of configurations. These parameters can additionally incorporate some level of slack to withstand an expected attack load in addition to an expected genuine workload.
“Normal” workload parameters:
M: max # of impression sites in an epoch contributing to non-zero loss in epoch.
N: max # of conversion sites that request non-zero loss from an epoch.
n: max # of conversion sites that request non-zero loss from a single (epoch, impression site) pair.
r: max budget consumed by an intermediary’s cross-advertiser queries on a single conversion site, as a fraction of the intermediary’s $\epsilon_{\textnormal{querier}}$.
Cap Capacity configuration
Per-querier filter $\epsilon_{\textnormal{querier}}$: configuration parameter
Global filter $\epsilon_{\textnormal{global}}=\max(N,n\cdot M)(1+r)\epsilon_{\textnormal{querier}}$
Impression-site quota $\epsilon_{\textnormal{imp-quota}}=n(1+r)\epsilon_{\textnormal{querier}}$
Conversion-site quota $\epsilon_{\textnormal{conv-quota}}=(1+r)\epsilon_{\textnormal{querier}}$
Quota count $k_{\textnormal{quota-count}}=2$ or 33
Tab. 2. Big Bird filter configurations.
4.7  Resilience guarantee
Quotas are configured so that each conversion and impression site can only enable consumption for a limited portion of the global budget. Therefore, an attacker’s impact on the global budget is proportional to the number of sites it controls. More precisely, since privacy loss only occurs when both impression sites and conversion sites are accessed (stock-and-flow pattern), an attacker needs to simultaneously control multiple impression sites and conversion sites to make a dent in the global budget. Any budget not consumed by the attacker can be used to serve genuine sites. Hence, Big Bird provides resilience to DoS depletion, where the attacker strength in Thm. 1 is expressed in number of attacker sites.
Theorem 1 (Resilience to DoS depletion – proof in C.3).
Consider an attacker who manages to create $M^{\textrm{adv}}$ and $N^{\textrm{adv}}$ imp-quota and conv-quota filters, respectively. The maximum budget $\epsilon_{\textnormal{global}}^{\textrm{adv}}$ that the attacker can consume from the global budget on a device $d$ is such that:
$\epsilon_{\textnormal{global}}^{\textrm{adv}}\leq\min(M^{\textrm{adv}}\epsilon_{\textnormal{imp-quota}},\ N^{\textrm{adv}}\epsilon_{\textnormal{conv-quota}}).$
It is easy to see how the preceding result implies that Big Bird structurally blocks the example attacks in Fig. 2. (a) The quota-count bound – small by design – severely constrains the number of quota budgets (in this case, conversion-site quota budgets) the attacker is able to leverage from a single user visit via, for example, automatic redirections. This prevents it from doing too much damage to the global budget. (b) There, all impressions and conversions occur under one domain, yielding $M^{\textrm{adv}}=N^{\textrm{adv}}=1$ and capping consumption at $\min(\epsilon_{\textnormal{imp-quota}},\epsilon_{\textnormal{conv-quota}})$, far from depletion.
While intuitive, Thm. 1 becomes meaningless when an attacker controls many sites (large $M^{\text{adv}}$ and $N^{\text{adv}}$). Domains names are cheap, but user actions are costly. This is why we provide an additional guarantee in Thm. 2, showing that the attacker’s damage is proportional to the number of user actions on a device. This result expands on Thm. 1 thanks to the quota count, which ties accessed impression and conversion sites to user actions. Indeed, an attacker who receives $U_{\textrm{adv}}$ interactions from user $u$ can create at most $M^{\textrm{adv}}+N^{\textrm{adv}}\leq k_{\textnormal{quota-count}}\cdot U^{\textrm{adv}}$. In other words, our quotas ensure graceful degradation for benign workloads as a function of $U^{\textrm{adv}}$. We prove the following:
Theorem 2 (Graceful degradation – proof here).
Consider an attacker collecting $U^{\textrm{adv}}$ user actions on sites under their control for device $d$. Under the configuration of Table 2, the budget $\epsilon_{\textnormal{global}}^{\textrm{adv}}$ that this attacker can consume from the global budget is upper-bounded by:
[…]
Proof.
[…]
Limitations. Thm. 1 and Thm. 2 show resilience of the global budget to DoS attacks, by bounding $\epsilon_{\textnormal{global}}^{\textrm{adv}}$. But can the quota budgets themselves become the target of DoS attacks? Unfortunately yes, although for limited forms of attacks that our threat model in §4.1 explicitly leaves out of scope. For instance, a conversion site’s quota can be depleted by malicious queriers, thereby preventing other intermediaries (or the conversion site itself) from running queries on that conversion site. Same for the impression site quota.
These attacks are less severe than attacks on the global budget for two reasons. First, damage is contained within attacked conversion or impression site and does not affect queriers operating on other sites. Second, the victims of these attacks (either conversion site or impression site) bear partial responsibility for the exposure – e.g., by displaying impressions from, or enabling measurement by, untrusted intermediaries. Finally, the quota configuration provides some level of resilience for quota budgets themselves: (1) for impression-site budgets, $n$ conversion sites are needed to deplete an impression site budget; (2) for conversion-site budgets, $(1+r)$ queriers are needed to deplete the budget. Still, we believe having top-level sites defend their quotas against malicious partners will require contractual agreements or other ecosystem-level measures such as registration, auditing, or partner vetting – beyond the reach of our budget-management layer, and thus outside our threat model.
4.8  Privacy guarantee
Our privacy analysis of Big Bird aims for high-fidelity modeling to provide confidence that our design satisfies global device–epoch–level IDP. The model, detailed in Appendix C (Algorithms 68), captures key aspects of Big Bird, including individual DP, on-device report generation with centralized noise addition, public information inherent to the protocol, data and budget adaptivity, multiple epochs and queriers, and the atomic interaction of multiple budgets within a transaction – making it one of the most faithful system models we have seen in the DP literature. At the time of this writing, the quota-count cap has not yet been incorporated into the model, though we plan to do so in a forthcoming revision and will update both the model, proof, and if necessary, the system’s design if the cap affects global device–epoch–level IDP. This process reflects a principle of our approach: to model systems as faithfully as possible while remaining explicit about known omissions and committed to revisiting the analysis as the design evolves.
Alg. 3 Formal model for Big Bird (simplified).
1:Input
2: Challenge bit $b\in\{0,1\}$
3: In-or-out record $x_{0}$ to be inserted at time $t_{0}$
4: Data generation process $\mathcal{G}$
5: Queriers $\mathcal{Q}^{1},\dots,\mathcal{Q}^{K}$
6: IDP budgets for each $x\in\mathcal{X}$: $\mathcal{F}_{x}^{\textnormal{global}}$, $(\mathcal{F}_{x}^{\textnormal{querier}}[k])_{k\in[K]}$, $(\mathcal{F}_{x}^{\textnormal{conv-quota}}[c])_{c\in[\mathcal{S}]}$, $(\mathcal{F}_{x}^{\textnormal{imp-quota}}[i])_{i\in[\mathcal{S}]}$
7:for $t\in[t_{\max}]$ do
8: // Data generation model:
[…]
10:if $t=t_{0}$ and $b=1$ then // insert in-our-out record on challenge
[…]
12: // Query execution model:
13:for $k\in[K]$ do
14: // Abstract query executed over the whole database
[…]
17:for $x\in D$ do
18: // Compute device-epoch loss and device-epoch-impSite loss (for imp-quota).
19:  $\epsilon_{x}\leftarrow\operatorname{EpochBudget}(x,q_{t}^{k})$
20:  for $i\in\mathbf{i}$ do
21:  $\epsilon_{x}^{i}\leftarrow\operatorname{EpochImpSiteBudget}(x,i,q_{t}^{k})$
22: // Atomic transaction across all budgets.
23:  if $(\mathcal{F}_{x}^{\textnormal{global}}\circ\mathcal{F}_{x}^{\textnormal{querier}}[k]\circ\mathcal{F}_{x}^{\textnormal{conv-quota}}[c]\circ\mathcal{F}_{x}^{\textnormal{imp-quota}}[\mathbf{i}])(\epsilon_{x},\epsilon_{x}^{\mathbb{i}})$ then
24: // Drop out-of-budget record.
25:  $S_{t}^{k}\leftarrow S_{t}^{k}\setminus\{x\}$
[…]
27:return $V^{1},\dots,V^{K}$
Alg. 3 shows a simplified version of our Appendix C model, but one in line with the generic model from §3.1. The purpose of including this simplified version in the body of the paper is to provide intuition on how we represent key aspects of Big Bird’s design. First, the budgets operate under IDP semantic is visible in the $x\in\mathcal{X}$ subscript at Line 6, where $x$ corresponds to a device-epoch. This indicates that the filters are each maintained on each device and for each epoch. Second, privacy loss is computed at a finer granularity (Line 21) for imp-quota than for the other budgets. Third, Big Bird uses a shared limit ($\mathcal{L}$ in §3) consisting of a collection of privacy and quotas budgets, all formally, privacy filters denoted $\mathcal{F}$. At Line 23, $(\mathcal{F}_{x}^{\textnormal{global}}\circ\dots\circ\mathcal{F}_{x}^{\textnormal{imp-quota}}[i])(\epsilon_{x},\epsilon_{x}^{i})$ denotes the atomic check and consumption across budgets, that happens for each record (device-epoch). For comparison, $\mathcal{F}_{x}^{\textnormal{global}}(\epsilon_{x})\wedge\dots\wedge\mathcal{F}_{x}^{\textnormal{imp-quota}}[i](\epsilon_{x}^{i})$ would check budgets independently, potentially spending from one but not the other, which harms utility. We show that this atomicity, never previously studied, preserves global IDP. Moreover, for the entire model of the system, we prove the following property:
Theorem 3 (Global IDP Guarantee – proof in §C.2).
Consider $x\in\mathcal{X}$ with global budget capacity $\epsilon_{\textnormal{global}}$. Then, Big Bird satisfies individual device-epoch $\epsilon_{\textnormal{global}}$-DP for $x$ under public information $\mathcal{C}$.
That is, for all data generation process $\mathcal{G}$, for all queriers $\mathcal{A}^{1},\dots,\mathcal{A}_{k}$, the output $V^{(b)}$ of Alg. 8 under challenge bit $b$ verifies:
(1) $\forall v\in\mathcal{R}^{e_{\max}\times(t_{\max}+1)\times B},\left|\ln\left(\frac{\Pr[V^{(0)}=v]}{\Pr[V^{(1)}=v]}\right)\right|\leq\epsilon_{\textnormal{global}}.$
4.9  Future: Resilience via fair scheduling
In an earlier draft of this arXiv paper, we proposed a scheduling algorithm to replace the impression-site quota budget, which can lead to under-utilization of the global budget. We omit that algorithm here, as further work is needed to reconcile its competing goals: improving global-budget utilization, matching the resilience guarantees of quotas, and preserving the global IDP guarantee. We summarize the direction briefly below, noting it is ongoing work in our group.
Static quotas risk under-utilizing the global budget: unused capacity at some sites is wasted while others are blocked, with impression-site quotas particularly problematic in this respect. To address this, we are exploring an on-device scheduling approach inspired by max–min fairness in operating systems, which provides isolation by guaranteeing each user a fair share while reallocating unused capacity to others. Our idea is to replace impression-site quotas with a short-interval scheduler that batches report requests and redistributes unused quota across impression sites, thereby improving utilization while retaining resilience to depletion. Unlike our earlier arXiv version, which assumed such scheduling would require delaying reports to queriers – something that Attribution editors considered undesirable – we now believe it can be implemented without report delays through adjustments to the underlying MPC/TEE protocol. The main open challenge is reconciling such scheduling with IDP semantics, since individual privacy loss cannot be used in scheduling decisions, which limits provable fairness and resilience guarantees.
5  Prototype
We implement Big Bird in two components: (1) pdslib, a general-purpose on-device individual DP library in Rust, and (2) its integration into Firefox’s Private Attribution, a basic Attribution prototype. pdslib: Big Bird’s core logic lives in pdslib, a Rust library for privacy budget management designed for broader individual DP use cases beyond advertising—e.g., location services in mobile apps. pdslib provides a generic interface: clients (sites or apps) register events (e.g., ad views, location visits) and request reports (e.g., attributions, model updates), receiving encrypted responses under strict privacy and isolation constraints. It implements all filters, quotas, privacy accounting, and cross-report optimizations. Big Bird is a Attribution-specific instantiation—a  350 LoC shim atop pdslib’s 2k LoC, specializing its generics to the Attribution spec.
Firefox integration: We integrate pdslib and the Big Bird shim into Firefox’s Private Attribution, replacing its primitive report-count-based accounting with full privacy loss tracking (Firefox’s PA lacks even Cookie Monster logic) (moz-pa, 12). Appendix E shows a Firefox extension dashboard we built to visualize filter and quota usage. We plan to open-source pdslib, the shim, and the integration to support PATWG and broader private-aggregation use cases.
6  Evaluation
We seek to answer the following questions:
  • (Q1)
    What parameters define “normal” operation in the Criteo workload?
  • (Q2)
    Do query error rates vary with different quota capacities?
  • (Q3)
    Do quotas preserve low error rates for benign queries under DoS attacks?
6.1  Methodology
Dataset. We evaluate Big Bird on CriteoPrivateAd (criteo2025, 40), a dataset released by the Criteo ad-tech company, a PATWG participant, for the purpose of benchmarking private advertising systems. The dataset samples 30 days of production traffic using third-party cookies, with 104M impressions across 220k publisher sites (publisher_id) and 10k conversion sites (campaign_id, a good proxy for advertiser domains (criteo2025, 40)). The data involves a single intermediary–Criteo itself.
Each impression includes contextual and user features, a daily-reset device ID, and attribution information indicating whether it led to a click, visit, or sale on a conversion site. This lets us reconstruct per-device, per-day conversion lists.
The dataset is impression-subsampled, not device-subsampled, so most devices have only one impression. Criteo provides the true device-level impression distribution and a resampling method to match it (criteo2025, 40). Using this, we construct a dataset with 4.6M impressions and 5.6M conversions across 1.4M devices, in which the median (resp. 90th percentile) device has 2 (resp. 6) impressions and 4 (resp. 16) conversions. We tune our algorithms and workload parameters on the first 10 days and report results on the remaining 20 days that we explicitly hold out for this evaluation.
Benign workload process. We consider a single-advertiser measurement scenario for benign queries. For each conversion, the advertiser–or Criteo acting on its behalf–invokes measureConversion and immediately after getReport() on the returned attributionObject, to request a report that attributes the conversion to the most recent relevant impression. Impressions are grouped into five buckets based on the features_ctx_not_constrained_0 field, which is anonymized but we assume it represents region, device type, or user category. The end-to-end DP query produces a per-advertiser histogram estimating the number of conversions attributed to each bucket. We adopt $\mathbf{RMSRE_{\tau}}$—relative root mean square error truncated at $\tau$—to measure DP histogram error against the ground-truth histogram, following (AksuGKKMSV24, 3).
When requesting attribution reports, each advertiser must specify a privacy budget, denoted by $\epsilon$. We assign this budget in a way that mimics how a real advertiser might choose it—by aiming for a certain level of accuracy in their reports. First, we determine how many conversions each advertiser typically sees per day. To ensure reports can be reasonably accurate under differential privacy, we only include advertisers that average at least 100 conversions daily. There are 73 such advertisers in the dataset. Next, we decide how many conversions to include in each aggregation batch. We set this batch size to be either ten times the advertiser’s daily average or 5,000 conversions—whichever is smaller. This ensures that advertisers produce roughly one report every 10 days, without aggregating on too small batches for low-volume advertisers. Finally, once we know how many conversions go into each batch, we choose the privacy budget $\epsilon$ so that the expected error in the report is about 5%, using a standard formula based on the Laplace mechanism and expected histogram statistics. This entire process is meant to reflect a realistic scenario, where advertisers select a privacy budget based on their volume and desired accuracy.
Attack workload process. To evaluate Big Bird’s resilience to DoS depletion attacks, we inject a synthetic adversarial workload into real benign traffic. Our setup simulates an attacker who controls 10 popular impression sites and 10 popular conversion sites, that each redirects to 7 new Sybil domains per real user action. This corresponds to a highly permissive configuration of quota-count =8=8 to ensure a strong attack, since most devices in Criteo convert only once. This attack evaluation methodology is still preliminary and we will explore other shapes of attacker traffic. We instantiate the attack as follows.
We first create 10 attacker impression sites, created by duplicating the most active impression sites to ensure the attacker interacts with many devices, since most devices in Criteo see only one impression site. For each impression, the attacker registers all its domains as target conversion sites. Then, we identify the top 10 real sites by number of conversions, and duplicate them to create 10 attacker conversion sites. For each user action on an attacker conversion site, we request a report with maximum $\epsilon=\epsilon_{\textnormal{querier}}$, where all the attacker impression sites are marked as relevant. We then redirect 7 times to new attacker conversion sites that request reports in the same way.
Since each series of attacker events (one impression or conversion event followed by 7 redirections) reuses device ids and timestamps from a real event, we need to break ties to decide in which order to inject attacker events on top of real events. We flip an unbiased coin to either let the attacker run first or the real event run first. This gives a fair chance for real sites to run their queries without being systematically front-runned by the attacker.
Baselines. We compare Big Bird against two baselines. The first is Attribution w/o global filter, which enforces only querier filters. This baseline is how Cookie Monster (TKM+24, 41) itself would behave. The second, Attribution w/ global filter, extends Attribution w/o global budget by adding a global budget, aligning with the current Attribution draft specification.
Defaults. Unless otherwise stated, we use $\epsilon_{\textnormal{querier}}=1$, $\epsilon_{\textnormal{conv-quota}}=1$, $\epsilon_{\textnormal{imp-quota}}=4$, and $\epsilon_{\textnormal{global}}=8$, reflecting our single-advertiser query workload (which implies $r=0$) and a filter configuration derived from a “normal” workload. We define this workload using the 95th percentile values of $N$, $M$, and $n$ shown in Tab. 3 and detailed in the next section. Finally, since Criteo resets user IDs daily, we fix epoch duration to 1 day.
(a) Error, normal workload
(b) Error causes, normal workload
(c) Error under attack
(d) Error causes under attack
Fig. 5. Quota system evaluation. (a), (b): Query error and its root causes in a benign case. (c), (d): Benign-query error and root causes under attack.
6.2  “Normal” workload parameters in Criteo (Q1)
Big Bird’s global and quota filter capacities are configured based on parameters intended to support a “normal” workload (Tab. 2). While our single-ad-tech dataset doesn’t provide reliable values for these parameters, we present a methodology that browser vendors can apply once Attribution is trialed at scale—and we illustrate this methodology on Criteo.
On a “training” dataset—the first 10 days of Criteo in our case—we can compute a distribution of $N,M,n$ values across devices, either by (1) running conversion attribution and computing their precise values as defined in Tab. 2, or (2) more efficiently, by computing upper bounds $\tilde{N}\geq N$, $\tilde{M}\geq M$, $\tilde{n}\geq n$ from the number of unique impression (resp. conversion) sites per device for $\tilde{M}$ (resp. $\tilde{N}$), and unique conversion sites per (device, impression-site) pair for $\tilde{n}$. We adopt the latter option for simplicity.
Tab. 3 shows these percentile values along with the corresponding global and impression-site quota capacities. Because our evaluation involves only single-advertiser queries, we force $r=0$ and thus $\epsilon_{\textnormal{conv-quota}}=\epsilon_{\textnormal{querier}}$. In more general settings, $r$ would also need to be set as a policy parameter.
[…]
p50 2 1 2 2 2
p90 4 2 4 8 4
p95 4 2 4 8 4
p99 6 3 6 18 6
p100 12 7 14 98 14
Tab. 3. Criteo “normal” workload.
Choosing quotas to support 100% of devices maximizes utility—since no quota-induced errors occur—but results in a very loose privacy budget, $\epsilon_{\textnormal{global}}=98$. A more balanced choice is the 95th percentile, which yields $\epsilon_{\textnormal{global}}=8$ and still avoids quota errors for the vast majority of devices. We validate the impact of this choice on accuracy in §6.3.
In general, selecting a percentile reflects a tradeoff between quota size (hence, utility) and the tightness of the global privacy guarantee $\epsilon_{\textnormal{global}}$. This tradeoff depends on workload characteristics: Criteo’s short epoch (1 day) and single-adtech scope suggest that real-world deployments—spanning multiple adtechs and longer epochs—will likely have higher $N,M,n$ values than in our table. For such workloads, stronger privacy guarantees may require adopting lower percentiles. We evaluate the effect of such tighter settings next.
6.3  Query errors under normal workload (Q2)
We vary $\epsilon_{\textnormal{imp-quota}}$ and measure its effect on query error in the benign case. Fig. 5(a) shows the median and tail (99th percentile) RMSRE. Since the Attribution baselines lack an imp-quota, their errors remains constant across $\epsilon_{\textnormal{imp-quota}}$ values. Moreover, they show identical error: at the p95 setting in Tab. 3, $\epsilon_{\textnormal{global}}=8$ is high enough to eliminate any error the global budget. In contrast, Big Bird’s error rises at low $\epsilon_{\textnormal{imp-quota}}$, as the quota forces some reports to be null, inducing error in query results. For $\epsilon_{\textnormal{imp-quota}}\geq 2$, the filter no longer affects query error, suggesting that reasonably sized quotas preserve utility. The p95 values from Tab. 3 are sufficient to support normal operation—and are even conservative, since actual privacy loss may be lower than the worst-case upper bounds we use to configure $N$, $M$, and $n$.
Fig. 5(b) breaks down the sources of Big Bird’s error. RMSRE stems from two factors: DP noise (variance), and null reports due to filter/quota blocking (bias). We isolate the latter by counting how many and which filters were out-of-budget for each report. If multiple filters are exceeded, we break ties in this order: querier, global, conv-quota, and imp-quota. We then compute the average fraction of affected reports per query. At $\epsilon_{\textnormal{imp-quota}}=1$, nearly a third of reports are blocked by imp-quota, explaining the high error in Fig. 5(a). For $\epsilon_{\textnormal{imp-quota}}\geq 2$, most blocked reports result from per-querier filters – which also exist in Attribution– explaining why Big Bird’s error converges to theirs.
6.4  Query errors under DoS attack (Q3)
We evaluate Big Bird’s resilience under attack against the global filter. Fig. 5(c) plots median and tail error for benign queries as a function of imp-quota capacity. Attribution w/o global filter is unaffected (but also lacks a global privacy guarantee). Attribution w/ global filter lacks protection from X, so its tail error rises sharply under attack. Big Bird with a well-configured quota (e.g., $\epsilon_{\textnormal{imp-quota}}=4$ per p95 in Tab. 3) fully isolates honest queriers, matching the error levels of Attribution w/o global filter.
As before, a too-small quota harms utility even without an attacker. But under attack, too-large quota ($\epsilon_{\textnormal{imp-quota}}\geq 7$) lets the attacker drain the global budget, degrading benign-query utility—consistent with our theoretical bounds. Fig. 5(d) confirms these observations by showing the break-down of error causes across imp-quota settings. At low $\epsilon_{\textnormal{imp-quota}}$, errors stem from honest reports blocked by the quota; the attack itself plays no role. At high $\epsilon_{\textnormal{imp-quota}}$, errors arise from global budget depletion, permitted by a loose imp-quota.
7  Related Work
Our main contribution advances the Attribution API—an emerging W3C standard poised to become the foundation for browser-based advertising measurement, and thus a critical part of the web’s infrastructure. Big Bird fills two key gaps in Attribution: it clarifies the semantics of per-site filters and introduces a system for configuring and managing both these filters and the global filter. This system upholds strong privacy guarantees, supports benign workloads, and resists global filter depletion. Big Bird also provides the basis of how budgeting should work for cross-advertiser queries, a planned Attribution extension.
Among prior work on Attribution and related APIs, the most relevant is Cookie Monster (TKM+24, 41), which tracks privacy loss using per-epoch filters tied to individual queriers. Big Bird goes further by managing these filters alongside the global filter and introducing a cross-report optimization that Cookie Monster lacks. Other foundational work includes Google’s ARA papers (AksuGKKMSV24, 3, 8, 15), research on the MPC components of these APIs (prio, 6, 7, 4), now-retired proposals like IPA (ipa, 20) and PAM (pam, 33), or Hybrid (hybrid-proposal, 19), which proposed several planned extensions to Attribution. More broadly, there is related work on privacy-preserving ad targeting (ibex, 44, 45).
Beyond ad measurement, this paper contributes to the broader challenge of privacy budget management, a crucial but understudied area in DP. While budget allocation within a single query is well-studied (matrix_mechanism, 28, 1), budget management across queries from mutually distrustful parties is less explored. Notably, (Pujol2021, 34, 35) balance utility across analysts sharing a single global budget, but without per-analyst guarantees. §3 additionally covers related work that neglected data adaptivity (GHH+25, 16, 41) and shared limits (DProvDB, 43, 24, 22, 14). Relevant systems include those for global budget scheduling (dpf, 30, 42, 23), which inspire our batched scheduling approach but differ in key ways. Most similar in spirit, (dpf, 30) proposes a max-min-fair algorithm for allocating global budget across epochs. Big Bird departs from these systems in two ways. First, it operates under epoch-level individual DP, which rules out relying on cross-epoch budget information. Second, it supports adaptive, multi-task queriers and defends against DoS depletion—gaps unaddressed in prior work.
Big Bird builds on the literature on privacy filters (filters_rogers, 36), particularly individual filters (FZ21, 13)—core DP primitives for adaptive composition and halting. Like prior work (aim, 31, 10, 27), we use filters to enforce DP. But we also repurpose them as quotas to limit consumption and provide isolation. A key contribution is our formalization of multi-granularity filter management—per-site and global—that supports simultaneous DP guarantees. These guarantees are essential in practice but have not been formalized correctly, as detailed in §3, nor has prior work shown how to manage them (DProvDB, 43, 24, 14). We do both.
8  Conclusions
This work revisited the foundations of privacy enforcement in the Attribution API and similar systems. We showed that per-querier budgeting, while practical and effective as an engineering constraint, has semantic limitations once data adaptivity and shared limits are taken into account. To preserve sound differential-privacy guarantees under these realistic conditions, we developed Big Bird, a global budget manager that embeds resilience to denial-of-service risks through a quota-based stock-and-flow model. Our analysis and evaluation demonstrate that rigorous privacy enforcement and robustness can coexist with utility, and that structured budget management can turn global differential privacy from a theoretical ideal into an operational mechanism suitable for deployment in standards like Attribution.
Appendix A Privacy Analysis for Per-Querier Guarantees
A.1  Data Adaptivity
We provide more details on the model introduced in 3.2. Our formalism with a data generation process follows existing work (LSV+19, 26). Other adaptive stream definitions have been proposed (GS13, 18, 21). Adaptive data generation definitions capture stronger adversaries than non-adaptive adjacency definitions for streams (CSS11, 5). Note that when there is only one querier, a single adversary can be data generation process and querier simultaneously.
Without loss of generality, the opt-out record and the time at which it is inserted are known upfront. We could let $\mathcal{G}$ adaptively decide when and what to insert, while enforcing that this happens at most once, but this decision would not actually depend on the challenge bit and would therefore follow a distribution known by $\mathcal{G}$ upfront.
Theorem 1.
Consider Alg. 1, where $\mathcal{F}^{k}$ are instantiated with pure DP filters (RRUV16, 37), each with capacity $\epsilon$. Denote by $V^{(b)}$ the output of the algorithm on challenge bit $b$.
When $K=1$, Alg. 1 reduces to single-querier fully adaptive composition on adaptively generated streams. We have:
$\forall x_{0}\in\mathcal{X},\forall t\in[t_{\max}],\forall\mathcal{G},\forall\mathcal{Q}^{1},$
$\forall S\subseteq\mathcal{R}^{t_{\max}},\left|\ln\left(\frac{\Pr[V^{1,(0)}\in S]}{\Pr[V^{1,(1)}\in S]}\right)\right|\leq\epsilon.$
However, for $K=n+1>1$, there exists $\epsilon^{\prime}>0$ such that:
$\exists x_{0}\in\mathcal{X},\exists t_{0}\in[t_{\max}],\exists\mathcal{G},\exists(\mathcal{Q}^{1},\dots,\mathcal{Q}^{n+1}),$
$\exists S\subseteq\mathcal{R}^{t_{\max}},\left|\ln\left(\frac{\Pr[V^{n+1,(0)}\in S]}{\Pr[V^{n+1,(1)}\in S]}\right)\right|=\epsilon+n\epsilon^{\prime}>\epsilon.$
Proof.
When $K=1$, we are exactly in the setting from (LSV+19, 26). Briefly, when data can change adaptively based on past results, inserting a single record can have cascading effects on future data and future queries. Still, after conditioning on a prefix $V_{1}^{1},\dots,V_{t}^{t}$, the databases in both worlds differ only by at most one record.
Now, consider $K=n+1$ with $n>0$. Take $x_{0}=1$ and $t_{0}=1$. Without loss of generality, take $t_{\max}=2$ and $\mathcal{X}=\{1\}$. We use histogram notation for the database, but since $\mathcal{X}$ is a singleton, we directly note $D\in\mathbb{N}$ instead of $D\in\mathbb{N}^{\mathcal{X}}$. In multiset notation, we could take $\sum_{x\in D}x$ to get an integer representation of the database. Define $\mathcal{G}$ and $\mathcal{Q}^{1},\dots,\mathcal{Q}^{n},\mathcal{Q}^{n+1}$ as follows:
  1. (1)
    At $t=1$:
    1. (a)
      The data generation process does not return any data: $\mathcal{G}(\emptyset,\dots,\emptyset)=\emptyset$. Thus $D_{1}=1$ if $b=1$, and $D_{1}=0$ otherwise, that is:
      (2) $D_{1}=b$
    2. (b)
      For $k\in\{1..n\}$, $\mathcal{Q}^{k}$ submits $q_{1}^{k}:D\mapsto D+\operatorname{Lap}(1/\epsilon)$, that consumes $\epsilon$ from the per-querier filter $\mathcal{F}^{k}$.
    3. (c)
      $\mathcal{Q}^{K}$ submits a query that consumes no budget from the per-querier filter: $\mathcal{Q}^{K}(\emptyset)=q_{1}^{K}:D\mapsto 0$.
  2. (2)
    At $t=2$:
    1. (a)
      The data generation process inserts up to $n$ new records, conditionally on outputs $V_{1}^{1},\dots,V_{1}^{n}$ from $\mathcal{Q}^{1},\dots,\mathcal{Q}^{n}$: $\mathcal{G}(V_{1}^{1},\dots,V_{1}^{K})=|\{k\in\{1..n\}:V_{1}^{k}>1/2\}|$. That is:
      (3) $D_{2}=\sum_{k=1}^{n}\mathds{1}[D_{1}+Z_{k}>1/2]$
      where $Z_{k}\sim\operatorname{Lap}(1/\epsilon)$. Since $V_{1}^{k}$ is more likely to be greater than $1/2$ if $b=1$, $D_{2}$ is more likely to be high if $b=1$. In other words, $\mathcal{G}$ encodes what $\mathcal{Q}^{1},\dots,\mathcal{Q}^{n}$ have learned into the dataset itself.
    2. (b)
      For $k\in\{1..n\}$, $\mathcal{Q}^{k}$ submits a query that consumes no budget from the per-querier filter: $\mathcal{Q}^{k}(V_{1}^{k})=q_{1}^{k}:D\mapsto 0$.
    3. (c)
      $\mathcal{Q}^{K}$submits a query that consumes $\epsilon$ from the per-querier filter $\mathcal{F}^{K}$: $\mathcal{Q}^{K}(V_{1}^{K})=q_{2}^{K}:D\mapsto D+\operatorname{Lap}(1/\epsilon)$, and therefore:
      (4) $V_{2}^{K}=D_{1}+D_{2}+Z$
      where $Z\sim\operatorname{Lap}(1/\epsilon)$. We will show that $V^{K}_{2}$ is disproportionally more likely to be high if $b=1$. In other words, $\mathcal{Q}^{K}$ retrieves the information encoded into the dataset, that originally came from $\mathcal{Q}^{1},\dots,\mathcal{Q}^{n}$.
Recall that $\Pr[Z_{k}>1/2]=\frac{1}{2}e^{-\epsilon/2}$. Also, $\Pr[1+Z_{k}>1/2]=\Pr[-Z_{k}<1/2]=1-\frac{1}{2}e^{-\epsilon/2}$, since $Z_{k}\sim Z_{k}$. We pose $p:=\frac{1}{2}e^{-\epsilon/2}$. $D_{2}$ follows a binomial distribution: $D_{2}^{(0)}\sim\mathcal{B}(n,p)$ and $D_{2}^{(1)}\sim\mathcal{B}(n,1-p)\sim n-\mathcal{B}(n,p)$. Hence, Eq. 4 gives:
(5) $\begin{cases}V_{2}^{K,(0)}\sim\mathcal{B}(n,p)+\operatorname{Lap}(1/\epsilon)\\ V_{2}^{K,(1)}\sim 1+n-\mathcal{B}(n,p)+\operatorname{Lap}(1/\epsilon)\end{cases}$
We already have $\mathbb{E}[V_{2}^{K,(0)}]=np<n/2$ and $\mathbb{E}[V_{2}^{K,(1)}]=1+n(1-p)>1+n/2$. We now exhibit an event for which $V^{K,(0)}_{2}\ll V^{K,(1)}_{2}$.
Take $B\sim\mathcal{B}(n,p)$. Take $N\geq n+1$. Recall that since $Z\sim\operatorname{Lap}(\epsilon)$, for all $z\geq 0$ and $t\in\mathbb{R}$ such that $z+t\geq 0$, we have:
(6) $\Pr[Z\geq z]=e^{\epsilon t}\Pr[Z\geq z+t]$
On one hand, we have $\Pr[V^{0}\geq N]=\Pr[B+Z\geq N]=\Pr[Z\geq N-B]$. On the other hand, $\Pr[V^{1}\geq N]=\Pr[1+n-B+Z\geq N]=\Pr[Z\geq N-1-n+B]$. Since $N-1-n+B\geq 0$, we have: $\Pr[Z\geq N-1-n+B]=e^{\epsilon}\Pr[Z\geq N-n+B]$. Hence,
(7) $\frac{\Pr[V^{1}\geq N]}{\Pr[V^{0}\geq N]}=e^{\epsilon}\frac{\Pr[Z\geq N-n+B]}{\Pr[Z\geq N-B]}$
For all $k\in[0,n]$ we have $N-n+k\geq 0$ and $N-k\geq 0$, so we can compute the exact probabilities:
$\Pr$ $[Z\geq N-n+B]=\sum_{k=0}^{n}{n\choose k}p^{k}(1-p)^{n-k}\Pr[Z\geq N-n+k]$
$=\sum_{k=0}^{n}{n\choose k}p^{k}(1-p)^{n-k}\frac{1}{2}e^{-(N-n+k)\epsilon}$
$=\frac{1}{2}e^{-N\epsilon}\sum_{k=0}^{n}{n\choose k}p^{k}(1-p)^{n-k}(e^{\epsilon})^{n-k}$
$=(p+(1-p)e^{\epsilon})^{n}$
Similarly,
$\Pr$ $[Z\geq N-B]=\sum_{k=0}^{n}{n\choose k}p^{k}(1-p)^{n-k}\Pr[Z\geq N-k]$
$=\frac{1}{2}e^{-N\epsilon}\sum_{k=0}^{n}{n\choose k}p^{k}(1-p)^{n-k}(e^{\epsilon})^{k}$
$=(pe^{\epsilon}+1-p)^{n}$
Hence,
$\frac{\Pr[V^{1}\geq N]}{\Pr[V^{0}\geq N]}=e^{\epsilon}\left(\frac{p+(1-p)e^{\epsilon}}{pe^{\epsilon}+1-p}\right)^{n}$
Moreover, we have:
$p+(1-p)e^{\epsilon}-(pe^{\epsilon}+1-p)$ $=(1-2p)e^{\epsilon}-(1-2p)$
$=(1-2p)(e^{\epsilon}-1)$
$>0$
since $p<1/2$ and $\epsilon>0$. Hence we can pose: $\epsilon^{\prime}:=\ln\left(\frac{p+(1-p)e^{\epsilon}}{pe^{\epsilon}+1-p}\right)>0$, which gives:
$\ln\left(\frac{\Pr[V^{1}\geq N]}{\Pr[V^{0}\geq N]}\right)=\epsilon+n\epsilon^{\prime}$
A.2  Shared Limits
Since we know that data adaptivity already threatens per-querier budget semantics, we consider a simpler setting where data is known upfront. Alg. 4 presents this variant of Alg. 1.
Alg. 4 Formal model for shared limits
1:Input
2: Database $D$
3: Queriers $\mathcal{Q}^{1},\dots,\mathcal{Q}^{K}$
4: Per-querier filters $\mathcal{F}^{1},\dots,\mathcal{F}^{K}$
5: Shared limit $\mathcal{L}$
6:Output
7: View $V=(V^{1},\dots,V^{K})\in\mathcal{R}^{t_{\max}\times K}$ of all $\mathcal{Q}^{k}$
8:for $t\in[t_{\max}]$ do
9:for $k\in[K]$ do
10: // Query based on $\mathcal{Q}^{k}$’s own previous results
[…]
12: // Execute if (1) $\mathcal{Q}^{k}$ has enough budget and (2) shared limit accepts the query.
13:if $\mathcal{F}^{k}(q_{1}^{k},\dots,q_{t}^{k})\wedge\mathcal{L}(q_{1}^{1},\dots,q_{1}^{K},\dots q_{t}^{k})$ then
14:  $V^{k}_{t}\leftarrow q_{t}^{k}(D)$
15:else
16:  $V^{k}_{t}\leftarrow\bot$
Theorem 2.
Take $K=n+1>1$, per-querier filters with capacity $\epsilon$ and shared filter with capacity $\epsilon_{g}<(n+1)\epsilon$ (allowing executions that trigger the shared limit).
For $n$ large enough to verify $\lceil\frac{n}{2}e^{-\epsilon/2}\rceil<n/2$, there exists neighboring databases $D,D^{\prime}$ such that:
$\exists(\mathcal{Q}^{1},\dots,\mathcal{Q}^{n+1}),\exists S\subseteq\mathcal{R}^{t_{\max}},$
$\left|\ln\left(\frac{\Pr[V^{n+1,(0)}\in S]}{\Pr[V^{n+1,(1)}\in S]}\right)\right|>\epsilon.$
Moreover, there exists $\epsilon^{\prime}>0$ and neighboring databases $D,D^{\prime}$ such that for all $n\geq\frac{4\ln 2}{(1-e^{-\epsilon/2})^{2}}$,
$\exists(\mathcal{Q}^{1},\dots,\mathcal{Q}^{n+1}),\exists S\subseteq\mathcal{R}^{t_{\max}},$
$\left|\ln\left(\frac{\Pr[V^{n+1,(0)}\in S]}{\Pr[V^{n+1,(1)}\in S]}\right)\right|\geq\epsilon+n\epsilon^{\prime}$
Proof.
Consider $D=0$, $D^{\prime}=1$ from $\mathcal{N}^{\mathcal{X}}=\mathbb{N}$ with $\mathcal{X}=\{1\}$. Define $\mathcal{Q}^{1},\dots,\mathcal{Q}^{n}$ as follows. Take $\epsilon_{0},\epsilon_{1},\epsilon_{2}>0$ that we will calibrate later.
  1. (1)
    At $t=1$:
    1. (a)
      For $k\in\{1..n\}$, $\mathcal{Q}^{k}$ submits $q_{1}^{k}:D\mapsto D+\operatorname{Lap}(1/\epsilon_{1})$, that consumes $\epsilon_{1}$.
    2. (b)
      Then, $\mathcal{Q}^{n+1}$ submits a query that consumes all but $\epsilon_{0}$ from its budget: $\mathcal{Q}^{n+1}(\emptyset)=q_{1}^{n+1}:D\mapsto D+\operatorname{Lap}(1/(\epsilon-\epsilon_{0}))$
  2. (2)
    At $t=2$:
    1. (a)
      For $k\in\{1..n\}$, $\mathcal{Q}^{k}$ submits a query that consumes no budget if $V_{1}^{k}$ was small, or $\epsilon_{2}$ if $V_{1}^{k}$ was large: $\mathcal{Q}^{k}(V_{1}^{k})=q_{2}^{k}:D\mapsto\mathds{1}[V_{1}^{k}>1/2](D+\operatorname{Lap}(1/\epsilon_{2}))$. In other words, $\mathcal{Q}^{k}$ encodes the information it learned at $t=1$ into the shared limit itself, through the adaptively chosen budget associated with its request at time $t=2$.
    2. (b)
      $\mathcal{Q}^{n+1}$ submits a query that attempts to consume $\epsilon_{0}$: $\mathcal{F}^{n+1}$: $\mathcal{Q}^{n+1}(V_{1}^{n+1})=q_{2}^{n+1}:D\mapsto D+\operatorname{Lap}(1/\epsilon_{0})$, and therefore $V_{2}^{n+1}\sim D+\operatorname{Lap}(1/\epsilon_{0})$ if the shared limit and the per-querier filter both accept the query, and $V_{2}^{n+1}=\bot$ otherwise. In other words, $\mathcal{Q}^{n+1}$ retrieves the information encoded into the shared limit, that originally came from $\mathcal{Q}^{1},\dots,\mathcal{Q}^{n}$’s own requests.
Now, we configure $\epsilon_{1}$ and $\epsilon_{2}$ to ensure that no query prior to $V_{2}^{n+1}$ gets filtered. That is, $\epsilon-\epsilon_{0}+n(\epsilon_{1}+\epsilon_{2})\leq\epsilon_{g}=n_{0}\epsilon$ and $\epsilon_{1}+\epsilon_{2}\leq\epsilon$. To saturate the global filter when the first $n$ queriers all ask for $\epsilon_{2}$, we want:
(8) $\epsilon-\epsilon_{0}+n(\epsilon_{1}+\epsilon_{2})=\epsilon_{g}$
This means that the state $f_{2}^{n,(b)}$ of the shared filter just before query $t=2,k=n+1$, on input $D^{(b)}$, is:
(9) $f_{2}^{n}=n\epsilon_{1}+\epsilon-\epsilon_{0}+\sum_{k=1}^{n}\Pr[V_{1}^{k}>1/2]\epsilon_{0}$
Pose $p=\frac{1}{2}e^{-\epsilon/2}$. We have:
(10) $\begin{cases}f_{2}^{n,(0)}\sim n\epsilon_{1}+\epsilon-\epsilon_{0}+\mathcal{B}(n,p)\epsilon_{2}\\ f_{2}^{n,(1)}\sim n\epsilon_{1}+\epsilon-\epsilon_{0}+(n-\mathcal{B}(n,p))\epsilon_{2}\end{cases}$
In particular, $\mathbb{E}[f_{2}^{n,(0)}]=\epsilon-\epsilon_{0}+n(\epsilon_{1}+p\epsilon_{2})$ and $\mathbb{E}[f_{2}^{n,(0)}]=\epsilon-\epsilon_{0}+n(\epsilon_{1}+(1-p)\epsilon_{2})$. We take $\epsilon_{2}$ so that $\mathcal{Q}^{n+1}$ is more likely to trigger the global filter on $D^{(1)}$ and more likely to succeed on $D^{(0)}$:
(11) $n\epsilon_{1}+\epsilon-\epsilon_{0}+n\epsilon_{2}/2+\epsilon_{0}=\epsilon_{g}$
Subtracting Eq. 8 from Eq. 11, we get $\epsilon_{0}-n\epsilon_{2}/2=0$, i.e., $\epsilon_{2}=2\epsilon_{0}/n$.
Define $B^{(b)}\sim b(n-B)+(1-b)B$, the number of queriers that ask for $\epsilon_{2}$ at $t=2$, with $B\sim\mathcal{B}(n,p)$ as in the proof for Thm. 1. By construction, $\epsilon_{0}$ passes $\mathcal{Q}^{n+1}$’s per-querier filter, so we can only return $\bot$ if $\epsilon_{0}$ hits the shared filter’s capacity:
(12) $V_{2}^{n+1,(b)}=\bot\iff f_{2}^{n,(0)}+\epsilon_{0}>\epsilon_{g}$
(13) $\iff n\epsilon_{1}+\epsilon-\epsilon_{0}+B^{(b)}\epsilon_{2}+\epsilon_{0}>\epsilon-\epsilon_{0}+n(\epsilon_{1}+\epsilon_{2})$
(14) $\iff B^{(b)}\epsilon_{2}+\epsilon_{0}>n\epsilon_{2}$
(15) $\iff 2B^{(b)}/n+1>2$
(16) $\iff B^{(b)}>n/2$
Now, let’s upper bound $\Pr[B^{(0)}>n/2]$ and lower bound $\Pr[B^{(1)}>n/2]$. By Hoeffding’s inequality, we have:
(17) $\Pr[B^{(0)}>n/2]$ $=\Pr[B>n/2]$
(18) $\leq\Pr[B\geq n/2]$
(19) $=\Pr[B-np\geq n(1/2-p)]$
(20) $\leq\exp{(-2(1/2-p)^{2}n)}$
Similarly:
(22) $\Pr[B^{(1)}>n/2]$ $=\Pr[n-B>n/2]$
(23) $=1-\Pr[B\geq n/2]$
(24) $\geq 1-\exp{(-2(1/2-p)^{2}n)}$
Finally, we pose $S=\{(v,\bot),v>1\}$. since we have $\Pr[V_{1}^{n+1,(1)}>1]=e^{\epsilon-\epsilon_{0}}\Pr[V_{0}^{n+1,(1)}>1]$, and that $V_{1}^{n+1,(b)}$ and $V_{2}^{n+1,(b)}$ are independent, we get:
(25) $\frac{\Pr[V^{n+1,(1)}\in S]}{\Pr[V^{n+1,(0)}\in S]}=e^{\epsilon-\epsilon_{0}}\frac{\Pr[V_{2}^{n+1,(1)}=\bot]}{\Pr[V_{2}^{n+1,(0)}=\bot]}$
(26) $=e^{\epsilon-\epsilon_{0}}\frac{\Pr[B^{(1)}>n/2]}{\Pr[B^{(0)}>n/2]}$
(27) $\geq e^{\epsilon-\epsilon_{0}}\left(\exp{(2(1/2-p)^{2}n)}-1\right)$
Now, since the median of $B$ is $\lfloor np\rfloor$ or $\lceil np\rceil$, we already have $\Pr[B^{(0)}>n/2]<1/2<\Pr[B^{(1)}>n/2]$ when $\lceil np\rceil<n/2$, which happens for $n$ large enough or $p$ small enough (i.e., $\epsilon$ large enough).
We can get a more informative bound asymptotically. Pose $\alpha_{n}:=\exp{((1/2-p)^{2}n)}$. We have $\frac{\Pr[V^{n+1,(1)}\in S]}{\Pr[V^{n+1,(0)}\in S]}\geq\alpha_{n}^{2}-1$. For $\alpha_{n}$ large enough, for instance any $\alpha_{n}\geq 2$, we have $\alpha_{n}^{2}-1\geq\alpha_{n}$. We have $\alpha_{n}\geq 2\iff n\geq\frac{\ln 2}{(1/2-p)^{2}}=\frac{4\ln 2}{(1-e^{-\epsilon/2})^{2}}$, so for such $n$ we get:
(28) $\frac{\Pr[V^{n+1,(1)}\in S]}{\Pr[V^{n+1,(0)}\in S]}$
(29) $\geq e^{\epsilon-\epsilon_{0}}\alpha_{n}$
(30) $=e^{\epsilon-\epsilon_{0}}(\exp{(1/2-p)^{2})})^{n}$
That is:
(31) $\left|\ln\left(\frac{\Pr[V^{n+1,(0)}\in S]}{\Pr[V^{n+1,(1)}\in S]}\right)\right|\geq\epsilon+n\epsilon^{\prime}>\epsilon$
with $\epsilon^{\prime}:=(1/2-p)^{2}>0$.
Remark about atomicity. We clarify notation for Alg. 1, Line 16, where a query is executed only if both the per-querier filter and the shared limit permit it. Taking $\mathcal{F}^{k}(q_{1}^{k},\dots,q_{t}^{k})\wedge\mathcal{L}(q_{1}^{1},\dots,q_{1}^{K},\dots q_{t}^{k})$ means that each filter takes its decision independently. In particular, if $\mathcal{F}^{k}$ accepts a query $q_{t}^{k}$ while $\mathcal{L}$ rejects it, $q_{t}^{k}$ will not be executed, but $\mathcal{F}^{k}$ will not know that fact when it is time to make a decision about $q_{t+1}^{k}$. That means $\mathcal{F}^{k}$ can be overly conservative, by deducting budget for queries that are in fact not executed. While this preserves privacy, it can hurt utility. We propose a solution to this conundrum in §4, thanks to a new form of atomic filter that is designed to avoid spending per-querier budget needlessly. This can be modeled by a new object $(\mathcal{F}^{k}\circ\mathcal{L})(q_{1}^{1},\dots,q_{1}^{K},\dots q_{t}^{k})$ that has access to all the previous queries. The shared state in this object means that per-querier budgets act as heuristics (as opposed to formal per-querier DP guarantees) unless a dedicated siloing assumption holds.
A.3  Extension to Individual DP filters
The model from Alg. 1 considers standard filters (RRUV16, 37), where queriers can directly observe whether a query gets rejected. However, systems such as Cookie Monster (TKM+24, 41) or ARA (GHH+25, 16) operate under Individual DP filters (FZ21, 13), which behave differently. Because IDP budget depends on the data itself, IDP filters cannot return $\bot$ explicitly when they are out of budget. Instead, they silently drop the out-of-budget individuals from the dataset, and return a noisy answer.
This raises the question: since filter outputs are not visible under IDP, do shared limits still leak extra information? We give a positive answer. First, we adapt the formalism to IDP filters in Alg. 5, with changes from Alg. 4 in blue. Next, Thm. 1 shows a variant of Thm. 2 for IDP filters. Finally, Thm. 1 immediately applies since the counterexample from the proof does not trigger any filter, and therefore works identically under IDP filters.
Alg. 5 Formal model for shared limits with IDP filters
1:Input
2: Database $D$
3: Queriers $\mathcal{Q}^{1},\dots,\mathcal{Q}^{K}$
4: Per-querier individual filters $(\mathcal{F}^{1}_{x},\dots,\mathcal{F}^{K}_{x})_{x\in D}$
5: Shared limit $(\mathcal{L}_{x})_{x\in D}$
6:Output
7: View $V=(V^{1},\dots,V^{K})\in\mathcal{R}^{t_{\max}\times K}$ of all $\mathcal{Q}^{k}$
8:for $t\in[t_{\max}]$ do
9:for $k\in[K]$ do
10: // Query based on $\mathcal{Q}^{k}$’s own previous results
[…]
12: // Execute if (1) $\mathcal{Q}^{k}$ has enough budget and (2) shared limit accepts the query.
[…]
14:for $x\in D$ do
15:  if $\mathcal{F}^{k}_{x}(q_{1}^{k},\dots,q_{t}^{k})\wedge\mathcal{L}_{x}(q_{1}^{1},\dots,q_{1}^{K},\dots q_{t}^{k})$ then
16: // Drop out-of-budget records.
17:  $S_{t}^{k}\leftarrow S_{t}^{k}\setminus\{x\}$
[…]
Theorem 1.
Take $K=n+1>1$, per-querier individual filters with capacity $\epsilon$ and shared individual filter with capacity $\epsilon_{g}<(n+1)\epsilon$ (allowing executions that trigger the shared limit).
There exists $\epsilon^{\prime}>0$ and neighboring databases $D,D^{\prime}$ such that
$\exists(\mathcal{Q}^{1},\dots,\mathcal{Q}^{n+1}),\exists S\subseteq\mathcal{R}^{t_{\max}},$
$\left|\ln\left(\frac{\Pr[V^{n+1,(0)}\in S]}{\Pr[V^{n+1,(1)}\in S]}\right)\right|\geq\epsilon+n\epsilon^{\prime}$
Proof.
We modify the execution from the proof for Thm. 2 to match IDP semantics, where records get dropped conditionally on past results. This emulates a form of data adaptivity, except that datasets $S_{2}^{n+1}$ now differ because of shared filters for various records, instead of through a data generation process. We then leverage the results from Thm. 1.
The changes in execution are as follows:
  • $D^{\prime}$ contains $n+1$ distinct records $\{x_{0},x_{1},\dots,x_{n}\}$ and $D$ only contains $\{x_{1},\dots,x_{n}\}$. We no longer have $D\in\mathbb{N}$, instead $D\in\mathbb{N}^{\mathcal{X}}$ with $\mathcal{X}=\{x_{0},\dots,x_{n}\}$.
  • At $t=1$, queriers submit $q_{1}^{k}:D\mapsto|\{x\in D:x=x_{0}\}|+\operatorname{Lap}(1/\epsilon)$. This consumes $x_{0}$’s full individual budget.
  • At $t=2$:
    • For $k\in[n]$, querier $k$ requests $x_{k}$ with full budget, conditionally on past results. That is:
      $\mathcal{Q}^{k}(V_{1}^{k})=q_{2}^{k}:D\mapsto\mathds{1}[V_{1}^{k}\leq 1/2])\cdot(|\{x\in D:x=x_{k}\}|+\operatorname{Lap}(1/\epsilon))$
    • Finally, querier $n+1$ requests all the records $x_{1},\dots,x_{n}$:
      $\mathcal{Q}^{n+1}(V_{1}^{n+1})=q_{2}^{n+1}:D\mapsto|\{x\in D:x\in\{x_{1},\dots,x_{n}\}\}|+\operatorname{Lap}(1/\epsilon)$
Since $\{x_{1},\dots,x_{n}\}$ are present in both worlds, $V^{{n+1},(1)}$ and $V^{{n+1},(0)}$ differ only by the points that get dropped at Line 17. Point $x_{k}$ gets dropped only if $V_{1}^{k}\leq 1/2$, which is more likely to happen on input $D$ (i.e., $b=0$). More precisely, we have:
(32) $V^{{n+1},(0)}\sim B+Z\text{ and }V^{{n+1},(1)}\sim n-B+Z$
with $B\sim\mathcal{B}(n,p)$ with $p=\frac{1}{2}\exp(-\epsilon/2)$ and $Z\sim\operatorname{Lap}(1/\epsilon)$.
This is exactly the same situation as Eq. 7, where the $e^{\epsilon}$ factor comes from the privacy loss at $t=1$. We conclude the same way.
A.4  Querier-Siloing Assumption
Lemma 0.
Consider Alg. 1 with $K\in\mathbb{N}$, per-querier filters with capacity $\epsilon>0$ and shared limit $\mathcal{L}$. Suppose that the following assumption holds:
(33) $\forall k\in[K],\forall u,w\in\mathcal{R}^{t}:u^{k}=w^{k},$
$\mathcal{L}\left(q_{1}^{1},\dots,q_{1}^{K},\dots,q_{t}^{1}(u^{1}_{<t}),\dots q_{t}^{k}(u^{k}_{<t})\right)\cdot q_{t}^{k}\left(D_{\leq t}(u_{<t})\right)$
$=\mathcal{L}\left(q_{1}^{1},\dots,q_{1}^{K},\dots,q_{t}^{1}(w^{1}_{<t}),\dots q_{t}^{k}(w^{k}_{<t})\right)\cdot q_{t}^{k}\left(D_{\leq t}(w_{<t})\right)$
Then, for all $k\in[K]$:
$\forall x_{0}\in\mathcal{X},\forall t_{0}\in[t_{\max}],\forall\mathcal{G},\forall(\mathcal{Q}^{1},\dots,\mathcal{Q}^{n+1}),$
$\forall S\subseteq\mathcal{R}^{t_{\max}},\left|\ln\left(\frac{\Pr[V^{k,(0)}\in S]}{\Pr[V^{k,(1)}\in S]}\right)\right|\leq\epsilon.$
In other words, if querier $k$’s results don’t depend on other queriers, either through data or shared limits, then Alg. 1 is per-querier $\epsilon$-DP for querier $k$.
Proof.
Take $x\in\mathcal{X}$, $\mathcal{G}$, $\mathcal{Q}^{1},\dots,\mathcal{Q}^{K}$, $k\in[K]$, and $v^{k}\in\mathcal{R}^{t}_{\max}$. We have:
$\Pr[V^{k}=v^{k}]$ $=\prod_{t=1}^{t_{\max}}\Pr[V_{t}^{k}=v_{t}^{k}|V_{1}^{k}=v_{1}^{k},\dots,V_{t-1}^{k}=v_{t-1}^{k}]$
Take $t\in[t_{\max}]$. We now express $V_{t}^{k}$ in terms of queries and filter decisions, since the filter properties are necessary for the privacy proof. Without loss of generality, we treat filters as functions that return 0 for reject and 1 for accept at Line 16, and we output 0 instead of $\bot$ at Line 19. Hence, we can express $V_{t}^{k}$ as:
(34) $V_{t}^{k}=\mathcal{F}^{k}(q_{1}^{k},\dots,q_{t}^{k})\cdot\mathcal{L}(q_{1}^{1},\dots,q_{1}^{K},\dots q_{t}^{k})\cdot q_{t}^{k}(D_{\leq t})$
The terms in this expression depend on various variables:
  • Once we condition on $v_{1}^{k},\dots,v_{t-1}^{k}$, we know the queries asked by $\mathcal{Q}^{k}$:
    (35) $q_{1}^{k}=\mathcal{Q}^{k}(),q_{2}^{k}=\mathcal{Q}^{k}(v_{1}^{k}),\dots,q_{t}^{k}=\mathcal{Q}^{k}(v_{1}^{k},\dots,v_{t-1}^{k}).$
  • As a consequence, $\mathcal{F}^{k}(q_{1}^{k},\dots,q_{t}^{k})$ is also fully determined by $v_{1}^{k},\dots,v_{t-1}^{k}$.
  • However, $\mathcal{L}(q_{1}^{1},\dots,q_{1}^{K},\dots q_{t}^{k})$ can depend on other queriers’ past results $V^{i}_{t}$ for $i\neq k$, since these outputs are used to select $q_{t}^{i}$.
  • Similarly, $D_{\leq t}$ can also depend on other queriers’ past results.
Pose $V^{\neq k}:=(V_{i}^{j})_{i\in[t],j\in[K]\setminus\{k\}}$. By total probabilities we have:
$\Pr[V_{t}^{k}=v_{t}^{k}|$ $V_{1}^{k}=v_{1}^{k},\dots,V_{t-1}^{k}=v_{t-1}^{k}]$
$=\int_{v^{\neq k}\in\mathcal{R}^{t\times(K-1)}}$ $\Pr[V_{t}^{k}=v_{t}^{k}|V_{1}^{k}=v_{1}^{k},\dots,V_{t-1}^{k}=v_{t-1}^{k},V^{\neq k}=v^{\neq k}]$
(36) $\cdot\Pr[V^{\neq k}=v^{\neq k}]dv^{\neq k}$
Take any $v^{\neq k}\in\mathcal{R}^{t\times(K-1)}$. Define $u\in\mathcal{R}^{t\times K}$ by concatenating $v^{k}$ and $v^{\neq k}$ such that $u^{k}=v^{k}$ and $u^{\neq k}=v^{\neq k}$. Thanks to Eq. 34 with explicit conditioning on all past results, we have:
$\Pr$ $[V_{t}^{k}=v_{t}^{k}|V_{1}^{k}=v_{1}^{k},\dots,V_{t-1}^{k}=v_{t-1}^{k},V^{\neq k}=v^{\neq k}]$
$=$ $~\mathcal{F}^{k}(q_{1}^{k},\dots,q_{t}^{k}(u^{k}_{<t}))$
$\cdot\mathcal{L}\left(q_{1}^{1},\dots,q_{1}^{K},\dots,q_{t}^{1}(u^{1}_{<t}),\dots q_{t}^{k}(u^{k}_{<t})\right)$
(37) $\cdot q_{t}^{k}\left(D_{\leq t}(u_{<t})\right)$
Define $\tilde{v}\in\mathcal{R}^{t\times K}$ by concatenating $v^{k}$ and $(\bot,\dots,\bot)\in\mathcal{R}^{t\times(K-1)}$. We have $\tilde{v}^{k}=u^{k}$, so by Eq. 33, Eq. 37 becomes:
(38) $\Pr$ $[V_{t}^{k}=v_{t}^{k}|V_{1}^{k}=v_{1}^{k},\dots,V_{t-1}^{k}=v_{t-1}^{k},V^{\neq k}=v^{\neq k}]$
(39) $=$ $~\mathcal{F}^{k}(q_{1}^{k},\dots,q_{t}^{k}(\tilde{v}^{k}_{<t}))$
(40) $\cdot\mathcal{L}\left(q_{1}^{1},\dots,q_{1}^{K},\dots,q_{t}^{1}(\tilde{v}^{1}_{<t}),\dots q_{t}^{k}(\tilde{v}^{k}_{<t})\right)$
(41) $\cdot q_{t}^{k}\left(D_{\leq t}(\tilde{v}_{<t})\right)$
This term is independent on $v^{\neq k}$, so we can express it as a function of $v^{k}_{<t}$ and take it out of the integral at Eq. 36, which gives:
(42) $\Pr$ $[V_{t}^{k}=v_{t}^{k}|V_{1}^{k}=v_{1}^{k},\dots,V_{t-1}^{k}=v_{t-1}^{k}]$
(43) $=\left(\mathcal{F}^{k}(q_{1}^{k},\dots,q_{t}^{k})\cdot\mathcal{L}(q_{1}^{1},\dots,q_{t}^{k})\cdot q_{t}^{k}(D_{\leq t})\right)\left(v^{k}_{<t}\right)$
Now we can bound the per-querier privacy loss:
$\left|\ln\left(\frac{\Pr[V^{k}=v^{k}]}{\Pr[V^{k}=v^{k}]}\right)\right|$
$=\left|\ln\left(\prod_{t=1}^{t_{\max}}\frac{\Pr[V_{t}^{k,(0)}=v_{t}^{k}|V_{1}^{k,(0)}=v_{1}^{k},\dots,V_{t-1}^{k,(0)}=v_{t-1}^{k}]}{\Pr[V_{t}^{k,(1)}=v_{t}^{k}|V_{1}^{k,(1)}=v_{1}^{k},\dots,V_{t-1}^{k,(1)}=v_{t-1}^{k}]}\right)\right|$
$\leq\sum_{t=1}^{t_{\max}}\left|\ln\left(\frac{\Pr[\mathcal{F}^{k}(q_{1}^{k},\dots,q_{t}^{k})\cdot\mathcal{L}(q_{1}^{1},\dots,q_{t}^{k})\cdot q_{t}^{k}(D_{\leq t}^{(0)})\{v^{k}_{<t}\}=v_{t}^{k}]}{\Pr[\mathcal{F}^{k}(q_{1}^{k},\dots,q_{t}^{k})\cdot\mathcal{L}(q_{1}^{1},\dots,q_{t}^{k})\cdot q_{t}^{k}(D_{\leq t}^{(1)})\{v^{k}_{<t}\}=v_{t}^{k}]}\right)\right|$
(44) $\leq\sum_{t=1}^{t_{\max}}\left|\ln\left(\frac{\Pr[\mathcal{F}^{k}(q_{1}^{k},\dots,q_{t}^{k})\cdot q_{t}^{k}(D_{\leq t}^{(0)})\{v^{k}_{<t}\}=v_{t}^{k}]}{\Pr[\mathcal{F}^{k}(q_{1}^{k},\dots,q_{t}^{k})\cdot q_{t}^{k}(D_{\leq t}^{(1)})\{v^{k}_{<t}\}=v_{t}^{k}]}\right)\right|$
The last inequality comes from what follows. $\mathcal{L}(q_{1}^{1},\dots,q_{t}^{k})$ is entirely determined by $v^{k}_{<t}$ and thus identical at the numerator and denominator. For $t$ where $\mathcal{L}(q_{1}^{1},\dots,q_{t}^{k})=1$, we have $\mathcal{F}^{k}(q_{1}^{k},\dots,q_{t}^{k})\cdot\mathcal{L}(q_{1}^{1},\dots,q_{t}^{k})\cdot q_{t}^{k}(D_{\leq t}^{(b)})=\mathcal{F}^{k}(q_{1}^{k},\dots,q_{t}^{k})\cdot q_{t}^{k}(D_{\leq t}^{(b)})$. For $t$ where $\mathcal{L}(q_{1}^{1},\dots,q_{t}^{k})=0$, we have:
$\left|\ln\left(\frac{\Pr[\mathcal{F}^{k}(q_{1}^{k},\dots,q_{t}^{k})\cdot\mathcal{L}(q_{1}^{1},\dots,q_{t}^{k})\cdot q_{t}^{k}(D_{\leq t}^{(0)})\{v^{k}_{<t}\}=v_{t}^{k}]}{\Pr[\mathcal{F}^{k}(q_{1}^{k},\dots,q_{t}^{k})\cdot\mathcal{L}(q_{1}^{1},\dots,q_{t}^{k})\cdot q_{t}^{k}(D_{\leq t}^{(1)})\{v^{k}_{<t}\}=v_{t}^{k}]}\right)\right|$
$=0\leq\left|\ln\left(\frac{\Pr[\mathcal{F}^{k}(q_{1}^{k},\dots,q_{t}^{k})\cdot q_{t}^{k}(D_{\leq t}^{(0)})\{v^{k}_{<t}\}=v_{t}^{k}]}{\Pr[\mathcal{F}^{k}(q_{1}^{k},\dots,q_{t}^{k})\cdot q_{t}^{k}(D_{\leq t}^{(1)})\{v^{k}_{<t}\}=v_{t}^{k}]}\right)\right|$
The rest of the proof is identical to standard filter guarantees on adaptively generated data. Suppose the filter uses an upper bound $\epsilon_{i}$ on the privacy loss for each query at step $i\in[t_{\max}]$:
(45) $\left|\ln\left(\frac{\Pr[q_{i}^{k}(D_{\leq i}^{(0)})\{v^{k}_{<i}\}=v_{i}^{k}]}{\Pr[q_{i}^{k}(D_{\leq i}^{(1)})\{v^{k}_{<i}\}=v_{i}^{k}]}\right)\right|\leq\epsilon_{i}(v^{k}_{<i})$
Then, by definition of pure DP filters, we get the following recursive property for $t\in[t_{\max}]$: $\mathcal{F}^{k}(q_{1}^{k},\dots,q_{t}^{k})=1$ iif $\sum_{i<t}\mathcal{F}^{k}(q_{1}^{k},\dots,q_{t}^{k})\cdot\epsilon_{i}(v^{k}_{<i})+\epsilon_{t}(v^{k}_{<t})\leq\epsilon$.
We conclude by applying this property at $t=t_{\max}$ in Eq. 44:
$\left|\ln\left(\frac{\Pr[V^{k}=v^{k}]}{\Pr[V^{k}=v^{k}]}\right)\right|\leq\epsilon$
Theorem 3.
Suppose that the two following assumptions hold:
  1. (1)
    Querier $k$’s queries operates on a restricted data stream that only depends on its own past results. That is, $D_{t}\leftarrow\mathcal{G}(V^{1}_{<t},\dots V^{K}_{<t})$ is of the following form:
    D_t^1 G(V^1_¡t), …, D_t^K G(V^K_¡t)
    with $q_{t}^{k}(D_{\leq t})=q_{t}^{k}(D_{\leq t}^{k})$.
  2. (2)
    Shared limits do not filter $k$’s reports depending on other queriers’ results:
    $\forall k\in[K],\forall u,w\in\mathcal{R}^{t}:u^{k}=w^{k},$
    $\mathcal{L}\left(q_{1}^{1},\dots,q_{1}^{K},\dots,q_{t}^{1}(u^{1}_{<t}),\dots q_{t}^{k}(u^{k}_{<t})\right)$
    $=\mathcal{L}\left(q_{1}^{1},\dots,q_{1}^{K},\dots,q_{t}^{1}(u^{1}_{<t}),\dots q_{t}^{k}(u^{k}_{<t})\right)$
    Sufficient conditions for this property to hold are if queriers are non-adaptive, never reach the shared limit, or operate on disjoint sets of devices (for on-device accounting with IDP filters).
Then the assumption from Lem. 3 holds, and thus per-querier guarantees hold.
Proof.
The first point implies:
$q_{t}^{k}(v_{<t}^{k})(D_{\leq t}(v_{<t}))=q_{t}^{k}(v_{<t}^{k})(D_{\leq t}(v_{<t}^{k}))$
Combining this with the second point, we directly get the assumption:
(46) $\forall k\in[K],\forall u,w\in\mathcal{R}^{t}:u^{k}=w^{k},$
$\mathcal{L}\left(q_{1}^{1},\dots,q_{1}^{K},\dots,q_{t}^{1}(u^{1}_{<t}),\dots q_{t}^{k}(u^{k}_{<t})\right)\cdot q_{t}^{k}\left(D_{\leq t}(u_{<t})\right)$
$=\mathcal{L}\left(q_{1}^{1},\dots,q_{1}^{K},\dots,q_{t}^{1}(w^{1}_{<t}),\dots q_{t}^{k}(w^{k}_{<t})\right)\cdot q_{t}^{k}\left(D_{\leq t}(w_{<t})\right)$
Appendix B Data and query model for Big Bird
This section sets up the formalism that underpins the main theorems, proved in subsequent sections and evoked in the body of the paper. We begin defining Big Bird’s data and query model, broadly adapted from Cookie Monster’s formalism. We then provide useful properties that will be needed in subsequent sections. Indeed, while Cookie Monster analyzes global and individual sensitivities of queries at the device-epoch level, Big Bird additionally needs such analyses at the device-epoch-site level.
Definition 0 (Database with per-site semantics).
A database $D$ is a set of device-epoch records where each record $x=(d,e,F)\in\mathcal{X}=\mathcal{D}\times\mathcal{E}\times\mathcal{P}(\mathcal{S}\times\mathcal{I}\cup\mathcal{S}\times\mathcal{P}(\mathcal{S})\times\mathcal{C})$ contains a device $d$, an epoch $e$ and a set of impression and conversion events $F$. Each event $f\in F$ contains the site (impression site $i$ or conversion site $c$) where the event occurred: $f=(i,\text{imp})\in\mathcal{S}\times\mathcal{I}$ or $f=(c,\mathbf{b},\text{conv})\in\mathcal{S}\times\mathcal{P}(\mathcal{S})\times\mathcal{C}$. Additionally, conversions contain the set of queriers $\mathbf{b}$ that will receive the conversion report ($\mathbf{b}=\{b\}$ without the cross-report privacy loss optimization from §D).
We now define queries and reports, with a slight adaptation of Cookie Monster’s definitions to our per-site semantics. While Cookie Monster comes with an arbitrary set of public events $P$ for each querier $b$, here for simplicity we assume that all the conversions for a querier are public. Using in Cookie Monster’s terminology, that means we set $P=\mathcal{C}_{b}:=\{(c,b,\text{conv}),c\in\mathcal{C},\text{conv}\in\mathcal{C}\}$, where $\mathcal{C}$ is the set of all conversions. Also, while Cookie Monster defines a set of relevant events, potentially including conversions, in Def. 2 we only consider relevant impressions for simplicity. In particular, this hardcodes ”Case 1” from (TKM+24, 41, Thm. 1).
Definition 0 (Attribution function, adapted from Cookie Monster).
Fix a set of relevant impression sites $\mathbf{i}_{A}\subset\mathcal{S}$ and a set of impressions relevant to the query $F_{A}\subset\mathbf{i}_{A}\times\mathcal{I}$. Fix $k,m\in\mathbb{N}^{*}$ where $k$ is a number of epochs. An attribution function is a function $A:\mathcal{P}(\mathcal{I})^{k}\to\mathbb{R}^{m}$ that takes $k$ event sets $F_{1},\dots,F_{k}$ from $k$ epochs and outputs an $m$-dimensional vector $A(F_{1},\dots,F_{k})$, such that only relevant events contribute to $A$. That is, for all $(F_{1},\dots,F_{k})\in\mathcal{P}(\mathcal{I})^{k}$, we have:
(47) $A(F_{1},\dots,F_{k})=A(F_{1}\cap F_{A},\dots,F_{k}\cap F_{A}).$
Definition 0 (Report identifier and attribution report, same as Cookie Monster).
Fix a domain of report identifiers $\mathbb{Z}$. Consider a mapping $d(\cdot)$ from report identifiers $R$ to devices $\mathcal{D}$ that gives the device $d_{r}$ that generated a report $r$.
Given an attribution function $A$, a set of epochs $E$ and a report identifier $r\in\mathbb{Z}$, the attribution report $\rho_{r,A,E}$, or $\rho_{r}$ for short, is a function over the whole database $D$ defined by:
(48) $\rho_{r}:D\in\mathbb{D}\mapsto A(D^{E}_{d_{r}}).$
Definition 0 (Query, same as Cookie Monster).
Consider a set of report identifiers $R\subset\mathbb{Z}$, and a set of attribution reports $(\rho_{r})_{r\in R}$ each with output in $\mathbb{R}^{m}$. The query for $(\rho_{r})_{r\in R}$ is the function $Q:\mathbb{D}\to\mathbb{R}^{m}$ is defined as $Q(D):=\sum_{r\in R}\rho_{r}(D)$ for $D\in\mathbb{D}$.
Definition 0 (Per-Epoch Sensitivity, same as (TKM+24, 41)).
Fix a report $\rho:\mathbb{D}\to\mathbb{R}^{m}$ for some $m$. We define the per-epoch individual $L_{1}$ sensitivity of $\rho$ for a device-epoch $x\in\mathcal{X}$ as follows:
(49) $\Delta(\rho):=\max_{D,D^{\prime}\in\mathbb{D}:\exists x\in\mathcal{X},D^{\prime}\sim_{x}D}\|\rho(D)-\rho(D^{\prime})\|_{1},$
where $D^{\prime}\sim_{x}D$ means $D^{\prime}$ and $D$ differ by the single record $x$.
We also define the per-epoch global $L_{1}$ sensitivity of $\rho$ as follows:
(50) $\Delta(\rho):=\max_{x\in\mathcal{X}}\Delta_{x}(\rho)$
Definition 0 (Per-Epoch-Site Sensitivity).
Fix a report $\rho:\mathbb{D}\to\mathbb{R}^{m}$ for some $m$, and an impression site $i\in\mathcal{S}$. We define the per-epoch-site individual $L_{1}$ sensitivity of $\rho$ for a device-epoch-site $x\in\mathcal{X},i\in\mathcal{S}$ as follows:
(51) $\Delta_{x,i}(\rho):=\max_{D,D^{\prime}\in\mathbb{D}:D^{\prime}\sim_{x,i}D}\|\rho(D)-\rho(D^{\prime})\|_{1}$
where $D^{\prime}\sim_{x,i}D$ means $D^{\prime}$ and $D$ differ on a single record $x$’s impressions on site $i$. That is, there exists $D_{0}\in\mathbb{D}$, a record $x\in\mathcal{X}$ such that $\{D,D^{\prime}\}=\{D_{0}+x,D_{0}+x^{i\to\emptyset}\}$, where $x^{i\to\emptyset}:=(d,e,F^{i\to\emptyset})$ is the record obtained by removing all the impressions on $i$ from $x$.
We also define the per-epoch-site global $L_{1}$ sensitivity of $\rho$ as follows:
(52) $\Delta_{i}(\rho):=\max_{x\in\mathcal{X}}\Delta_{i,x}(\rho)$
To simplify subsequent results, we define some notation:
Definition 0 (Zeroing-out).
Fix a vector of impression sets $\mathbf{F}=(F_{1},\dots,F_{k})\in\mathcal{P}(\mathcal{I})^{k}$ for any $k>0$. For $i\in\mathcal{S}$ and $j\in[k]$ we define:
  • $\mathbf{F}^{j\rightarrow\emptyset}:=(F_{1},\dots,F_{j-1},\emptyset,F_{j+1},\dots,F_{k})$, i.e., we zero-out the $j$th epoch.
  • $\mathbf{F}^{j,i\rightarrow\emptyset}:=(F_{1},\dots,F_{j-1},F_{j}\setminus\mathcal{I}_{i},F_{j+1},\dots,F_{k})$, i.e., we zero-out all the impressions $\mathcal{I}_{i}:=\{(i,\text{imp}),\text{imp}\in\mathcal{I}\}$ belonging to site $i$ from the $j$th epoch.
Plugging Def. 2 into Def. 5 and Def. 6 immediately gives:
Lemma 0 (Global sensitivity of reports).
Fix a device $d$, a set of $k$ epochs $E$, an attribution function $A$ and the corresponding report $\rho:D\mapsto A(D^{E}_{d})$. We have:
(53) $\Delta(\rho)=\max_{\begin{subarray}{c}\mathbf{F}\in\mathcal{P}(\mathcal{I})^{k},j\in[k]\end{subarray}}\|A(\mathbf{F})-A(\mathbf{F}^{j\to\emptyset})\|_{1}.$
Moreover, for any impression site $i\in\mathcal{S}$ we have:
(54) $\Delta_{i}(\rho)=\max_{\begin{subarray}{c}\mathbf{F}\in\mathcal{P}(\mathcal{I})^{k},j\in[k],i\in\mathcal{S}\end{subarray}}\|A(\mathbf{F})-A(\mathbf{F}^{j,i\to\emptyset})\|_{1}.$
Theorem 9 (Individual sensitivity of reports per epoch-site).
Fix a report identifier $r$, a device $d_{r}$, a set of epochs $E_{r}=\{e_{1}^{(r)},\ldots,e_{k}^{(r)}\}$, a set of sites $I_{r}=\{i_{1},\ldots,i_{m_{e}}\}$ for each epoch $e\in E_{r}$, an attribution function $A$ with relevant events $F_{A}$, and the corresponding report $\rho:D\mapsto A(D^{E_{r},\{I_{r}\}_{e\in E_{r}}}_{d_{r}})$. Fix a device-epoch record $x=(d,e,F_{j})\in\mathcal{X}$, where $F\subseteq\mathcal{S}\times\mathcal{S}\times\mathcal{C}\cup\mathcal{S}\times\mathcal{I}$, so that $x_{i}=(d,e,F_{j,i})$ is the projection where $F_{i}$ contains only events related to site $i$.
We can upper bound the individual sensitivity of reports per epoch-site $\Delta_{x,i}(\rho)$ by:
[…]
Proof.
Fix a report $\rho$, an impression site $i$ and $x=(d,e,F)\in\mathcal{X}$ with impressions $F_{i}$ on site $i$. Consider any $D,D^{\prime}\in\mathbb{D}$ such that $D^{\prime}=D+x_{i}$. We have $\rho(D)=A(D_{d_{r}}^{E_{r},\{I_{r}\}_{e\in E_{r}}})$ and $\rho(D^{\prime})=A((D^{\prime})_{d_{r}}^{E_{r},\{I_{r}\}_{e\in E_{r}}})$.
  • First, if $d\neq d_{r}$, $e\notin E_{r}$, or $i\notin I_{r}$, then $(D^{\prime})_{d_{r}}^{E_{r},\{I_{r}\}_{e\in E_{r}}}=D_{d_{r}}^{E_{r},\{I_{r}\}_{e\in E_{r}}}$. Hence, $\|\rho(D)-\rho(D^{\prime})\|_{1}=0$ for all such $D,D^{\prime}$, which implies $\Delta_{x_{i}}(\rho)=0$.
  • Next, suppose that the report requests a single epoch $E_{r}=\{e_{r}\}$ with a single site $I^{(e_{r})}_{r}=\{i_{r}\}$:
    • If $d=d_{r}$, $e=e_{r}$, and $i=i_{r}$, then since $D+x_{i}=D^{\prime}$, we must have $(d_{r},e_{r},F_{i})\notin D$, and thus $D_{d_{r}}^{e_{r},i_{r}}=\emptyset$. On the other hand, $(D^{\prime})_{d_{r}}^{e_{r},i_{r}}=F_{i}$ (restricted to events relevant to site $i_{r}$). Thus, $\|\rho(D)-\rho(D^{\prime})\|_{1}=\|A(F_{i})-A(\emptyset)\|_{1}$.
    • If $d\neq d_{r}$, $e\neq e_{r}$, or $i\neq i_{r}$, then $(d,e,F_{i})$ doesn’t change the outcome and $(D^{\prime})_{e_{r}}^{i_{r}}=D_{e_{r}}^{i_{r}}$. Hence, $\|\rho(D)-\rho(D^{\prime})\|_{1}=0$.
  • Now, suppose that the report requests either an arbitrary range of epochs $E_{r}$ each of whom has at least one site, or a single epoch that has multiple sites $I^{(e_{r})}_{r}$:
    • If $d\neq d_{r}$, $e\notin E_{r}$, or $i\notin I_{r}^{(e)}$, then $A((D^{\prime})_{d_{r}}^{E_{r},\{I^{(e)}_{r}\}_{e\in E_{r}}})=A(D_{d_{r}}^{E_{r},\{I^{(e)}_{r}\}_{e\in E_{r}}})$, i.e., $\|\rho(D^{\prime})-\rho(D)\|_{1}=0$.
    • If we have $d=d_{r}$, $e=e_{j}^{(r)}\in E_{r}$, and $i\in I_{r}^{(e)}$, but $F_{i}$ is simply not related to the attribution request, i.e. $F_{i}\cap F_{A}=\emptyset$. Then, by definition of $F_{A}$, we have $A((D^{\prime})_{d_{r}}^{E_{r},\{I^{(e)}_{r}\}_{e\in E_{r}}})=A(D_{d_{r}}^{E_{r},\{I^{(e)}_{r}\}_{e\in E_{r}}})$, i.e., $\|\rho(D)-\rho(D^{\prime})\|_{1}=0$.
    • Otherwise, it must be the case that $d=d_{r}$, $e=e_{j}^{(r)}\in E_{r}$, $i\in I^{(e)}_{r}$ and $F_{i}\cap F_{A}\neq\emptyset$ and there are events in the intersection that is related to some site $i$ in epoch $e$, so we have:
      (56) $\|\rho(D)-\rho(D^{\prime})\|_{1}=\|A(\mathbf{F}^{j,i\to\emptyset})-A(\mathbf{F})\|_{1},$
      where $j$ is the index of epoch $e$ in $E_{r}$, and $F_{j,i}$ represents the relevant events for site $i$ in epoch $e_{j}^{(r)}$.
    The first two cases are independent over choices of $D\sim D^{\prime}$, so taking the max over such choices still gives $\Delta_{x_{i}}(\rho)=0$. Unfortunately, the third identity does depend on the choice of $D\sim D^{\prime}$, and taking the max only gives the general definition of global sensitivity, in the worst case.
Next, we show that $\Delta_{i}(\rho)\leq 2\Delta(\rho)$ for any report. For reports using certain attribution functions, we can have a tighter $\Delta_{i}(\rho)=\Delta(\rho)$ bound, but it does not hold in general.
Lemma 0 (Relationship between Device-Epoch-Site and Device-Epoch Sensitivities).
For any report $\rho$ with attribution function $A$, let $F$ denote the full dataset, $F_{j}$ denote all data in $F$ pertaining to epoch $j$, and $F_{j,i}$ denote all data in $F_{j}$ pertaining to site $i$. The following inequality holds:
(57) $\Delta_{i}(\rho)\leq 2\Delta(\rho)$
Proof.
Recall the definition of device-epoch-site global sensitivity:
(58) $\Delta_{i}(\rho)=\max_{\begin{subarray}{c}\mathbf{F}\in\mathcal{P}(\mathcal{I})^{k},j\in[k],i\in\mathcal{S}\end{subarray}}\|A(\mathbf{F}^{j,i\to\emptyset})-A(\mathbf{F})\|_{1}.$
We can decompose this expression using the triangle inequality:
(59) $\|A(\mathbf{F}^{j,i\to\emptyset})-A(\mathbf{F})\|_{1}$
(60) $=$ $\|A(\mathbf{F}^{j,i\to\emptyset})-A(\mathbf{F}^{j\to\emptyset})+A(\mathbf{F}^{j\to\emptyset})-A(\mathbf{F})\|_{1}$
(61) $\leq$ $\|A(\mathbf{F}^{j,i\to\emptyset})-A(\mathbf{F}^{j\to\emptyset})\|_{1}+\|A(\mathbf{F}^{j\to\emptyset})-A(\mathbf{F})\|_{1}$
For the first term, note that $A(\mathbf{F}^{j,i\to\emptyset})$ uses events $F_{j,\neg i}$ from epoch $j$ where $F_{j,\neg i}=F_{j}\setminus F_{j,i}$, while $A(\mathbf{F}^{j\to\emptyset})$ uses no events from epoch $j$. Since $F_{j,\neg i}\subseteq F_{j}$ and all events from epoch $j$ for a given device come from a single device-epoch record, we can view this as the change from adding a single record containing events $F_{j,\neg i}$. This is bounded by the definition of $\Delta(\rho)$:
(62) $\|A(\mathbf{F}^{j,i\to\emptyset})-A(\mathbf{F}^{j\to\emptyset})\|_{1}\leq\Delta(\rho).$
The second term represents the sensitivity to removing events from the entire epoch $j$. In the individual DP setting with device-epoch records, all events from epoch $j$ for a given device come from a single device-epoch record. Removing this entire record corresponds exactly to one of the cases considered in the definition of $\Delta(\rho)$. Therefore:
(63) $\|A(\mathbf{F}^{j\to\emptyset})-A(\mathbf{F})\|_{1}\leq\Delta(\rho).$
Therefore, substituting the two upper bounds into Eq. (61):
(64) $\Delta_{i}(\rho)\leq\Delta(\rho)+\Delta(\rho)=2\Delta(\rho).$
Appendix C Global Filter Management
C.1  Algorithm
Overview. Big Bird manages per-site and global privacy filters using the quota mechanisms described in §4.5. Alg. 6 depicts the functionality triggered on receiving a report request (i.e., measureConversion() and getReport()). Big Bird checks and consumes budget from the relevant filters and prunes the resulting report based on filter status. First, Big Bird ensures all filters and quotas are initialized. Second, it computes the privacy losses incurred—both at the epoch level (TKM+24, 41, §C) and at the epoch-impression-site level (Def. 1). Third, it checks whether all filters have sufficient budget and attempts to consume it. To ensure avoid wasting budget from some filters (or quotas) when other filters are out of budget, Big Bird uses a two-phase commit protocol to deduct privacy losses from multiple filters atomically. Alg. 7 formalizes this: if any filter cannot afford its share of the privacy loss, the report is zeroed out, and no budget is consumed from any filter for that report.
Subroutines. Def. 2, 1, 3 and Alg. 7 define subroutines used in Alg. 6. The first two definitions rely on the sensitivity bounds from Thm. 9 (for the impression-site quota) and (TKM+24, 41) (for the filters and the other quotas).
Definition 0 (EpochImpSiteBudget).
Let $x=(d,e,F)\in\mathcal{X}$ be a device-epoch record, $i\in S$ an impression site, $\rho$ an attribution report, and $\lambda>0$ the Laplace noise scale applied to $\rho$. Given the upper bound $\tilde{\Delta}_{x_{i}}(\rho)\geq\Delta_{x_{i}}(\rho)$ on the per-epoch-site individual sensitivity given by Thm. 9, we can upper bound the epoch–site privacy loss consumed by $\rho$ at $(x,i)$ by
(65) $\text{EpochImpSiteBudget}(x,i,\rho,\lambda):=\frac{\tilde{\Delta}_{x_{i}}(\rho)}{\lambda}.$
Definition 0 (EpochBudget, from (TKM+24, 41)).
Fix a device-epoch record $x\in\mathcal{X}$, $\rho$ an attribution report, and $\lambda>0$ the Laplace noise scale applied to $\rho$. Given the upper bound $\tilde{\Delta}_{x}(\rho)\geq\Delta_{x}(\rho)$ on the per-epoch individual sensitivity given by (TKM+24, 41), the individual privacy loss for device-epoch record $x$ is:
(66) $\operatorname{EpochBudget}(x,\rho,\lambda):=\frac{\tilde{\Delta}_{x}(\rho)}{\lambda}$
Definition 0 (Filters $\mathcal{F}_{x}$).
For each device-epoch record $x=(d,e,F)$, we maintain several pure DP filters (RRUV16, 37):
  • $\mathcal{F}^{\textnormal{querier}[b]}_{x}$ for each querier $b$,
  • $\mathcal{F}^{\textnormal{global}}_{x}$,
  • $\mathcal{F}^{\textnormal{conv-quota}[c]}_{x}$ for each conversion site $c$,
  • $\mathcal{F}^{\textnormal{imp-quota}[i]}_{x}$ for each impression site $i$.
We define $\operatorname{InitializeFilters}(\epsilon_{\textnormal{global}},\epsilon_{\textnormal{querier}},\epsilon_{\textnormal{imp-quota}},\epsilon_{\textnormal{conv-quota}})$ the function that returns these filters instantiated with respective capacities $\epsilon_{\textnormal{global}},\epsilon_{\textnormal{querier}},\epsilon_{\textnormal{imp-quota}}$, $\epsilon_{\textnormal{conv-quota}}$.
For each filter $\mathcal{F}$, we adapt the notation from  (RRUV16, 37) to recursively define $\mathcal{F}.\operatorname{canConsume}$, $\mathcal{F}.\operatorname{tryConsume}$ and $\text{pass}_{\mathcal{F}}[t]$, for a sequence of adaptively chosen privacy budgets $\epsilon_{x}^{1},\dots,\epsilon_{x}^{t}$, as follows:
  • $\operatorname{canConsume}(\epsilon^{t}_{x})$: Returns TRUE if the filter can accommodate additional privacy loss $\epsilon^{t}_{x}$, i.e., $\epsilon^{t}_{x}\leq\epsilon_{\text{initial}}-\sum_{k\in[t-1]}\epsilon^{k}_{x}\cdot\text{pass}_{\mathcal{F}}[k]$ where $\epsilon_{\text{initial}}$ is the filter capacity.
  • $\operatorname{tryConsume}(\epsilon^{t}_{x})$: Calls $\operatorname{canConsume}(\epsilon^{t}_{x})$. If successful, sets $\text{pass}_{\mathcal{F}}[t]=1$ to deduct $\epsilon$ from the filter’s remaining capacity; otherwise, sets $\text{pass}_{\mathcal{F}}[t]=0$.
Alg. 6 Big Bird algorithm (on-device)
1:Input
2: Budget capacities $\epsilon_{\textnormal{global}},\epsilon_{\textnormal{querier}},\epsilon_{\textnormal{imp-quota}},\epsilon_{\textnormal{conv-quota}}$
3:// Generate report and update on-device budget
4:def $\operatorname{GenerateReport}$($D,\rho,\lambda$)
5: Read $\rho$ to get device $d$, conversion site $c$, querier $k$, impression sites $\mathbf{i}$, target epochs $E$, attribution function $A$.
6:for $e\in E$ do
[…]
8:if $\mathcal{F}_{x}$ is not defined then
9: // Set up per-device filters and quotas (Def. 3)
10:  $\mathcal{F}_{x}\leftarrow\operatorname{InitializeFilters(\epsilon_{\textnormal{global}},\epsilon_{\textnormal{querier}},\epsilon_{\textnormal{imp-quota}},\epsilon_{\textnormal{conv-quota}})}$
[…]
14:for $i\in\mathbf{i}$ do
15:  $\epsilon_{x}^{i}\leftarrow\operatorname{EpochImpSiteBudget}(x,i,\rho,\lambda)$ // Def. 1
16:  $\epsilon_{\mathbf{x}}^{\mathbf{i},t}[i]\leftarrow\epsilon_{x}^{i}$
17: // Check and update filters/quotas atomically (Alg. 7)
18:if $\operatorname{AtomicFilterCheckAndConsume}(\mathcal{F}_{x},k,c,\mathbf{i},\epsilon^{t}_{x},\epsilon_{\mathbf{x}}^{\mathbf{i},t})=\textsf{FALSE}$ then
19:  $F_{e}\leftarrow\emptyset$ // Empty the epoch if any check fails
[…]
21:return $\rho$
Alg. 7 AtomicFilterCheckAndConsume
1:Input
2: Budgets $\mathcal{F}_{x}$, querier $b$, conversion site $c$, impression sites $\mathbf{i}$
[…]
5:// Phase 1: check if all filters can consume (Def. 3)
6:if $\mathcal{F}_{x}^{\textnormal{querier}[b]}.\operatorname{canConsume}(\epsilon^{t}_{x})=\textsf{FALSE}$ then
7:return FALSE
8:if $\mathcal{F}^{\textnormal{global}}_{x}.\operatorname{canConsume}(\epsilon^{t}_{x})=\textsf{FALSE}$ then
9:return FALSE
10:if $\mathcal{F}^{\textnormal{conv-quota}[c]}_{x}.\operatorname{canConsume}(\epsilon^{t}_{x})=\textsf{FALSE}$ then
11:return FALSE
12:for $i\in\mathbf{i}$ do
13:if $\mathcal{F}^{\textnormal{imp-quota}[i]}_{x}.\operatorname{canConsume}(\epsilon_{\mathbf{x}}^{\mathbf{i},t}[i])=\textsf{FALSE}$ then
14:return FALSE
15:// Phase 2: consume from all filters (Def. 3)
[…]
19:for $i\in\mathbf{i}$ do
[…]
21:return TRUE
C.2  Global Privacy Guarantees
Execution model. Alg. 8 presents an abstract model of Big Bird’s operation, capturing how it answers queries sequentially for all queriers. At each time step $t$, querier $k$ can formulate a query $Q_{t}^{k}$, which is a batch of reports as defined in Def. 4. When a querier has nothing to ask, it can send an empty query with zero sensitivity. AnswerQuery then processes each individual report $p_{r}$ (for $r\in R$, where $R$ is the set of report identifiers in $Q_{t}^{k}$). For each report, GenerateReport creates an individual clipped attribution report. Finally, reports are summed and noise before returning the query result.
Alg. 8 Formalism for DP analysis
1:Input
2: Challenge bit $b\in\{0,1\}$
3: Opt-out device $x_{0}=(d_{0},e_{0},F_{0})$
4: Data generation process $\mathcal{G}$
5: Queriers $\mathcal{Q}^{1},\dots,\mathcal{Q}^{K}$
6:Output
7: View $V=(V^{1},\dots,V^{B})\in\mathcal{R}^{e_{\max}\times(t_{\max}+1)\times K}$ of all $\mathcal{Q}^{k}$
8:for $e\in[e_{\max}]$ do
9:// Generation happens based on all previous results
[…]
12:if $e=e_{0}$ and $(d_{0},e_{0})\not\in G$ then
13: // Insert opt-out record, with or without private events.
14:if $b=0$ then
15:  $D^{e}\leftarrow D^{e}+(d_{0},e_{0},F_{0}\cap\mathcal{C})$
16:else
17:  $D^{e}\leftarrow D^{e}+(d_{0},e_{0},F_{0})$
18:// Each querier receives its conversions and corresponding report identifiers
[…]
20:for $(d,e,F)\in D_{e}$ do
21:for $f\in F:f=(c,\mathbf{b},\text{conv})$ do
22:  Generate report identifier $r\overset{{\scriptscriptstyle\mathdollar}}{\leftarrow}U(\mathbb{Z})$
23:// Save mapping from $r$ to the device that generated it
24:  $d_{r}\leftarrow d$
25:  for $k\in[K]$ do
26:  $S_{k}\leftarrow S_{k}\cup\{(r,f)\}$
27:for $k\in[K]$ do
28:  $V^{k}_{e,0}\leftarrow S_{k}$
29:// Queriers ask queries, round robin
30:for $t\in[t_{\max}]$ do
31:for $k\in[K]$ do
32: // Querier $k$ choses based on its own previous results
33:  $Q_{t}^{k},\lambda_{t}^{k}\leftarrow\mathcal{Q}^{k}(V^{k}_{\leq e,<t})$
34:  $V^{k}_{e,t}\leftarrow\operatorname{AnswerQuery}(D^{\leq e},Q_{t}^{k},\lambda_{t}^{k},k)$
35:// Collect, aggregate and noise reports to answer $Q$
36:def $\operatorname{AnswerQuery}$($D,Q,\lambda$)
[…]
38:for $r\in R$ do
[…]
40: Sample $X\sim\operatorname{Lap}(\lambda)$
41:return $\sum_{r\in R}\hat{\rho}_{r}+X$
Theorem 3.
Consider $x\in\mathcal{X}$ with global budget capacity $\epsilon_{\textnormal{global}}$. Then, Big Bird satisfies individual device-epoch $\epsilon_{\textnormal{global}}$-DP for $x$ under public information $\mathcal{C}$. That is, for all data generation process $\mathcal{G}$, for all queriers $\mathcal{A}^{1},\dots,\mathcal{A}_{k}$, the output $V^{(b)}$ of Alg. 8 under challenge bit $b$ verifies:
(67) $\forall v\in\mathcal{R}^{e_{\max}\times(t_{\max}+1)\times B},\left|\ln\left(\frac{\Pr[V^{(0)}=v]}{\Pr[V^{(1)}=v]}\right)\right|\leq\epsilon_{\textnormal{global}}.$
Proof.
Take a device-epoch $x=(d,e,F)\in\mathcal{X}$. Denote by $x_{\mathcal{C}}=(d,e,F\cap\mathcal{C})$ the device-epoch obtained by keeping only public events $\mathcal{C}$ from $x$, where public events are the set of all conversions. Take $v\in\mathcal{R}^{e_{\max}\times(t_{\max}+1)\times K}$ and $b\in\{0,1\}$. $v$ is the vector of values returned at different points in Alg. 8. We split it into $v=(v_{1,0},v_{1,1},\ldots,v_{1,t_{\max}},\dots,v_{e,0},\dots v_{e,t})$ where $v_{e,0}=(v_{e,0}^{1},\dots,v_{e,0}^{K})$ is a value for the output from Line 28 of Alg. 8 (public events from epoch $e$ for each querier $k\in[K]$), and $v_{e,t}=(v_{e,t}^{1},\dots,v_{e,t}^{K})$, where $v_{e,t}^{k}$ is the output for the query $Q^{k}_{t}$ for querier $k$ at step $t$ from epoch $e$ (Line 34). By conditioning over past outputs at each time step $(e,t,k)\in[e_{\max}]\times\{0,...,t_{\max}\}\times[K]$ we get:
(68) $\Pr[V^{(b)}=v]$
(69) $=$ $\prod_{e=1}^{e_{\max}}\prod_{t=0}^{t_{\max}}\prod_{k=1}^{K}\Pr[V^{k,(b)}_{e,t}=v_{e,t}^{k}|v_{\leq e,<t}^{<k}].$
Take $e\in[e_{\max}]$, $t\in[t_{\max}]$ and $k\in[K]$. By Algorithm 8, we have:
(70) $\Pr[V^{k,(b)}_{e,t}=v_{e,t}^{k}|v_{\leq e,<t}^{<k}]$
(71) $=$ $\Pr[\operatorname{AnswerQuery}(D^{(b)},Q^{k}_{t},\lambda_{t}^{k};\mathcal{F}^{(b)}_{e,t})=v_{e,t}^{k}],$
where the database $D^{(b)}$, the query $Q^{k}_{e,t}$ with associated noise $\lambda_{t}^{k}$ and the state of the privacy filters $\mathcal{F}^{k,(b)}_{e,t}$ are functions of past results $v_{\leq e,<t}^{<k}$. Queries are chosen by $\mathcal{Q}^{k}$ conditionally on past results but independently of $b$. Additionally, the database can depend on $b$ since we insert private data for the opt-out record conditionally on that bit at Line 14. The state of the privacy filters can also depend on $b$, albeit in a limited fashion, since we have:
(72) $\forall x^{\prime}\in D^{(0)}\cap D^{(1)},(\mathcal{F}^{k,(0)}_{e,t})_{x^{\prime}}=(\mathcal{F}^{k,(1)}_{e,t})_{x^{\prime}}$
Indeed, Alg. 6 shows that the budget spent for $x^{\prime}$ at each step only depends on past queries, which are identical in both worlds once we condition on $v_{\leq e,<t}^{<k}$, and on the record $x^{\prime}$ itself, which is identical in both worlds. This is thanks to device-epoch level accounting in Def. 2,1.
Now, if we denote by $\rho_{r}(D;\mathcal{F})=\operatorname{GenerateReport}(D,\rho_{r},\lambda;\mathcal{F})$ the filtered report returned by Alg. 6 (which maintains the state $\mathcal{F}$), we get:
(73) $\Pr[V^{k,(b)}_{e,t}=v_{e,t}^{k}|v_{\leq e,<t}^{<k}]=\Pr\left[\sum_{r\in R_{e,t}^{k}}\rho_{r}(D^{(b)};\mathcal{F}_{e,t,r}^{(b)})+X_{e,t}=v_{e,t}^{k}\right],$
where $X_{e,t}$ is the Laplace noise added at time $e,t$.
We omit the $e,t,k$ indices when they are clear from context. We denote by $D$ the database generated conditionally on $v_{\leq e,<t}^{<k}$ before adding the opt-out record, giving: $D^{(0)}=D+x_{\mathcal{C}}$ and $D^{(1)}=D+x$. By Def. 2, since $F_{A}\subset\mathcal{I}$ and $F\cap\mathcal{C}\cap\mathcal{I}=\emptyset$, we have:
$\rho_{r}(D^{(0)};\mathcal{F}_{r}^{(0)})=\rho_{r}(D+(d,e,F\cap\mathcal{C}\cap F_{a})=\rho_{r}(D;\mathcal{F}_{r}^{(0)})$
By Alg. 6, remark that the filter $(\mathcal{F}_{r}^{(0)})_{x^{\prime}}$ for $x^{\prime}\in\{x,x_{\mathcal{C}}\}$ has no impact on the output: replacing the corresponding data by $\emptyset$ at Line 19 is a no-op. Combining this observation with Eq. 72 gives:
(74) $\rho_{r}(D^{(0)};\mathcal{F}_{r}^{(0)})=\rho_{r}(D;\mathcal{F}_{r}^{(1)})$
We can now bound the privacy loss at step $e,t,b$:
(75) $\left|\ln\left(\frac{\Pr[V^{k,(0)}_{e,t}=v_{e,t}^{k}|v_{\leq e,<t}^{<k}]}{\Pr[V^{k,(1)}_{e,t}=v_{e,t}^{k}|v_{\leq e,<t}^{<k}]}\right)\right|$
(76) $=\left|\ln\left(\frac{\Pr\left[\sum_{r\in R_{e,t}^{b}}\rho_{r}(D;\mathcal{F}_{r}^{(1)})+X_{e,t}=v_{e,t}^{k}\right]}{\Pr\left[\sum_{r\in R_{e,t}^{b}}\rho_{r}(D+x;\mathcal{F}_{r}^{(1)})+X_{e,t}=v_{e,t}^{k}\right]}\right)\right|$
We now omit $\mathcal{F}_{r}^{(1)}$ since it is present on both sides. Take $r\in R_{t}$. Denote by $\text{pass}_{r}$ the output of Alg. 7 saying whether $x$ passes the atomic check for report $x$ in the world where $x$ is present. If $\text{pass}_{r}=0$, we have $\rho_{r}(D+x)=\rho_{r}(D)$ because of Alg. 6, Line 19. Otherwise, if $\text{pass}_{r}=1$, we have $\left\|\rho_{r}(D+x)-\rho_{r}(D)\right\|_{1}\leq\Delta_{x}(\rho_{r})$. Hence,
(77) $\|\rho_{r}(D)-\rho_{r}(D+x)\|\leq\Delta_{x}(\rho_{r})\text{pass}_{r}$
Thus by triangle inequality followed by Def. 2 we have:
(78) $\|\sum_{r\in R_{t}}\rho_{r}(D)-\sum_{r\in R_{t}}\rho_{r}(D+x)\|$ $\leq\sum_{r\in R_{t}}\Delta_{x}(\rho_{r})\text{pass}_{r}$
(79) $\leq\sum_{r\in R_{t}}\lambda_{t}^{k}\cdot\epsilon_{r}\text{pass}_{r}$
And since $X_{t}\sim\operatorname{Lap}(\lambda_{t}^{k})$, by property of the Laplace distribution Eq. 76 becomes:
(80) $\left|\ln\left(\frac{\Pr[V^{k,(0)}_{e,t}=v_{e,t}^{k}|v_{\leq e,<t}^{<k}]}{\Pr[V^{k,(1)}_{e,t}=v_{e,t}^{k}|v_{\leq e,<t}^{<k}]}\right)\right|\leq\sum_{r\in R_{t}}\epsilon_{r}\text{pass}_{r}$
We can now bound the total privacy loss. First, we use Eq. 68 and the fact that the outputs $v_{e,0}^{b}$ at Line 28 are identical across both worlds by definition of $x_{\mathcal{C}}$ to drop indices where $t=0$:
(81) $\left|\ln\left(\frac{\Pr[V^{(0)}=v]}{\Pr[V^{(1)}=v]}\right)\right|$
(82) $=\left|\ln\left(\prod_{e=1}^{e_{\max}}\prod_{t=0}^{t_{\max}}\prod_{b=1}^{B}\frac{\Pr[V^{k,(0)}_{e,t}=v_{e,t}^{k}|v_{\leq e,<t}^{<k}]}{\Pr[V^{k,(1)}_{e,t}=v_{e,t}^{k}|v_{\leq e,<t}^{<k}]}\right)\right|$
(83) $=\left|\sum_{e=1}^{e_{\max}}\sum_{t=1}^{t_{\max}}\sum_{b=1}^{B}\ln\left(\frac{\Pr[V^{k,(0)}_{e,t}=v_{e,t}^{k}|v_{\leq e,<t}^{<k}]}{\Pr[V^{k,(1)}_{e,t}=v_{e,t}^{k}|v_{\leq e,<t}^{<k}]}\right)\right|$
By triangle inequality and Eq. 80, we get:
(84) $\left|\ln\left(\frac{\Pr[V^{(0)}=v]}{\Pr[V^{(1)}=v]}\right)\right|$
(85) $\leq\sum_{e=1}^{e_{\max}}\sum_{t=1}^{t_{\max}}\sum_{b=1}^{B}\sum_{r\in R_{t}}\epsilon_{r}\text{pass}_{r}$
Finally, since $\text{pass}_{r}=1$ implies that $r$ passes $\mathcal{F}^{\textnormal{global budget}}$, by definition of $\mathcal{F}^{\textnormal{global budget}}$, the accumulated loss over all reports related to query $x$ is below the filter capacity:
(86) $\sum_{e=1}^{e_{\max}}\sum_{t=1}^{t_{\max}}\sum_{b=1}^{B}\sum_{r\in R_{t}}\epsilon_{r}\text{pass}_{r}\leq\epsilon_{\textnormal{global}}$
Hence we have shown Eq. 67. ∎
C.3  DoS Resilience
This section proves our main resilience result for Big Bird’s quota-based online algorithm: Thm. 1. First, Lem.  4 shows that the 2-PC check (Alg.  7) ensures (1) atomic consumption across all filters relevant to a query, and (2) when all filters have sufficient budget, each consumes an amount proportional to its level-specific sensitivity—either at epoch or at epoch-site level. Next, Lem.  5 and 7 bound the total privacy budget the adversary can consume from the global budget at any qualified time, using the atomic consumption guarantees of Lem. 4. This final bound directly implies Thm. 1.
We first formalize the atomicity property of the 2-PC algorithm for consuming privacy budgets from relevant filters when Big Bird answers a query at any time step $k$ (Alg. 7).
Lemma 0 (2-phase commit filter guarantees).
For any individual report generation request (e.g., , for a report $r$ processed by $\operatorname{GenerateReport}$ when invoked by $\operatorname{AnswerQuery}$ for a batch query $Q^{t}$ in Alg. 8, or more generally, any call to $\operatorname{AtomicFilterCheckAndConsume}$ in Alg. 6 and Alg. 7), let:
(87) $\text{pass}(r)=\begin{cases}1&\text{if $\operatorname{AtomicFilterCheckAndConsume}$ returns}\\ &\hskip 14.22636pt\text{{TRUE}\ for report $r$}\\ 0&\text{otherwise}\end{cases}$
If $\text{pass}(r)=1$, then $\operatorname{AtomicFilterCheckAndConsume}$ guarantees the following properties for that specific report $r$: The $\operatorname{AtomicFilterCheckAndConsume}$ function in Alg. 7 guarantees the following properties:
For any query $k$ processed by $\operatorname{AtomicFilterCheckAndConsume}$, if $\text{pass}(k)=1$, then
  1. (1)
    Epoch-level Consistency Property: exactly the same amount of budget $\epsilon^{t}_{x}$ is consumed by the querier budget, global budget, and conversion-site quota-filter for that query.
  2. (2)
    Epoch-site-level Consistency Property: exactly $\epsilon^{i}_{x}[i]$ is consumed by the impression-site quota filter, which represents the device-epoch-impressionsiteimpressionsite-level individual privacy loss.
Proof.
We can prove both properties at the same time. Fix an arbitrary individual report request, let’s denote it by $k$ for consistency within this proof, for which $\text{pass}(k)=1$ (as defined in the lemma statement, meaning $\operatorname{AtomicFilterCheckAndConsume}$ returns TRUE for this report $k$). From Alg. 7, we observe that $\operatorname{AtomicFilterCheckAndConsume}$ returns TRUE for this individual report $k$ if and only if: (1) all $\operatorname{canConsume}$ checks in Phase 1 pass, and (2) All $\operatorname{tryConsume}$ operations in Phase 2 are executed. For this individual report $k$, originating from conversion site $c_{k}$ with querier $b_{k}$ and intended impression sites $\mathbf{i}_{k}$, the function calls:
  • $\mathcal{F}_{\textnormal{querier budget}}[b_{k}].\operatorname{tryConsume}(\epsilon_{x}^{\text{report }k}$)
  • $\mathcal{F}_{\textnormal{global budget}}.\operatorname{tryConsume}(\epsilon_{x}^{\text{report }k}$)
  • $\mathcal{F}_{\text{conv-quota}}[c_{k}].\operatorname{tryConsume}(\epsilon_{x}^{\text{report }k}$)
  • For each $i\in\mathbf{i}_{k}$: $\mathcal{F}_{\text{imp-quota}}[i].\operatorname{tryConsume}(\epsilon_{x}^{\text{report }k}[i]$)
Note that $\epsilon_{x}^{\text{report }k}$ is the device-epoch-level individual privacy loss computed for this specific report $k$ (e.g., , via $\operatorname{EpochBudget}$ in Line 12 of Alg. 6). Similarly, each $\epsilon_{x}^{\text{report }k}[i]$ is the device-epoch-site-level individual privacy loss for impression site $i$ relevant to this report $k$ (via $\operatorname{EpochImpSiteBudget}$ in Alg. 6). Therefore, when $\text{pass}(k)=1$, the conversion-site quota filter (for $c_{k}$), the per-site filter (for $b_{k}$), and the global budget all consume exactly the same amount $\epsilon_{x}^{\text{report }k}$. Concurrently, each relevant impression-site quota filter (for $i\in\mathbf{i}_{k}$) consumes its specific amount $\epsilon_{x}^{\text{report }k}[i]$, which is proportional to its sensitivity at the epoch-impression-site level for this report $k$. ∎
With such atomic guarantees for every individual report processed up to some batch query step $t$ (as per Alg. 8), we can show upper bounds on how much an adversary can deplete the global budget by the end of batch query step $t$, on any device-epoch $x\in\mathcal{X}$. The main isolation result is in Thm. 1, which uses Lem. 56 and 7.
Notation. First, we introduce some notation describing adversarial behavior, which we use in subsequent proofs:
  • At step $t$ in Line 30, suppose that querier $b$ requests a report $\rho_{r,E,A}$ with noise $\lambda$ through conversion site $c$ for impression sites $\mathbf{i}$. Consider a device-epoch $x$, with individual budget $\epsilon^{t}_{x}$ computed at Line 12 in Alg. 6.
  • A report at time step $k$ concerns with one conversion site $c_{k}$, and a set of impression sites $\mathbf{i}_{k}\subseteq S$.
  • The union of attackers can control an arbitrary subset of conversion sites $\mathbf{bad_{c}}\subseteq S$. We denote by $N^{\mathsf{adv}}$ the size of $|\mathbf{bad_{c}}|$ over the entire lifetime. Similarly, the adversary can control an arbitrary subset of impression sites $\mathbf{bad_{i}}\subseteq S$. We denote by $M^{\mathsf{adv}}$ the size of $|\mathbf{bad_{i}}|$ over the entire lifetime. We let $\mathbf{bad}=\mathbf{bad_{c}}\cup\mathbf{bad_{i}}$ and $\mathbf{good}=S\backslash\mathbf{bad}$.
  • Denote by $N^{\leq t,\mathsf{adv}}$ the number of conversion sites in $\mathbf{bad_{c}}$ with respect to $x$ that were queried with non-zero budget by step $t$. Denote by $M^{\leq t,\mathsf{adv}}$ the number of impression sites in $\mathbf{bad_{i}}$ with respect to $x$ that were queried with non-zero budget by step $t$.
Lemma 0.
Consider a sequence of T batch query steps. If an adversary, across all individual report generation attempts within these T steps, successfully leads to the instantiation (via $\operatorname{saveImpression}$ calls) and subsequent use (successful budget consumption from) of at most $M^{\text{adv}}$ distinct imp-quota filters and $N^{\text{adv}}$ distinct conv-quota filters in an attempt to deplete the global budget, then the adversary consumes at most $M^{\mathsf{adv}}\epsilon_{\textnormal{imp-quota}}$ budget from the global budget.
Proof.
The lemma considers a sequence of $T$ batch query steps. Let $s$ be the index for these batch query steps, from $1$ to $T$. The total privacy loss in the global budget incurred by adversarial report generations up to (not including) batch query step $T$ (i.e., after T1T-1 steps have completed) is:
(88) $\epsilon_{\text{global}}^{\leq T-1,\mathbf{bad}}=\sum_{s=1}^{T-1}\sum_{r\in R_{s}:c_{r}\in\text{bad}_{c}\land\text{pass}(r,s)=1}\epsilon_{r,s}^{\text{global}}.$
Here, $s$ indexes batch query steps, $R_{s}$ is the set of individual reports in batch $s$, $c_{r}$ is the conversion site for report $r$, $\text{bad}_{c}$ are adversarial conversion sites, $\text{pass}(r,s)=1$ indicates report $r$ in batch $s$ was successfully processed, and $\epsilon_{r,s}^{\text{global}}$ is the global budget portion consumed by that individual report $r$ in batch $s$. By the consistency property of Lem. 4, for each successfully processed adversarial report $r$ associated with a conversion site $c_{r}\in\text{bad}_{c}$, the filter $\mathcal{F}_{\text{conv-quota}}[c_{r}]$ precisely tracks the privacy loss $\epsilon_{r,s}^{\text{global}}$. Thus, we can write:
(89) $\mathcal{F}_{\text{conv-quota}}[c_{r}]^{\leq T-1}=\sum_{s=1}^{T-1}\sum_{r^{\prime}\in R_{s}:c_{r^{\prime}}=c_{r}\land\text{pass}(r^{\prime},s)=1}\epsilon_{r^{\prime},s}^{\text{global}}.$
The quantity $\epsilon_{\textnormal{conv-quota}}[c_{r}]^{\leq T-1}$ (defined as $\mathcal{F}_{\text{conv-quota}}[c_{r}]^{\leq T-1}$ in the preceding equation) represents the sum of all $\epsilon_{r^{\prime},s}^{\text{global}}$ terms for reports $r^{\prime}$ associated with a specific conversion site $c_{r}$ up to step T1T-1. Therefore, by summing $\epsilon_{\textnormal{conv-quota}}[c]^{\leq T-1}$ over all adversarial conversion sites $c\in\mathbf{bad_{c}}$, we are effectively re-summing all the individual $\epsilon_{r,s}^{\text{global}}$ contributions that constitute $\epsilon_{\text{global}}^{\leq T-1,\mathbf{bad}}$. Thus, it follows directly from the definitions and equation (88) that:
(90) $\epsilon^{\leq T-1,\mathbf{bad}}_{\mathsf{global}}=\sum_{c\in\mathbf{bad_{c}}}\epsilon_{\textnormal{conv-quota}}[c]^{\leq T-1}.$
This sum can be restricted to conversion sites with non-zero privacy loss, i.e.:
(91) $=\sum_{c\in\mathbf{bad_{c}}:\epsilon_{\textnormal{conv-quota}}[c]^{\leq T-1}>0}\epsilon_{\textnormal{conv-quota}}[c]^{\leq T-1}$
(92) $\leq\sum_{c\in\mathbf{bad_{c}}:\epsilon_{\textnormal{conv-quota}}[c]^{\leq T-1}>0}\epsilon_{\textnormal{conv-quota}},$
where $\epsilon_{\textnormal{conv-quota}}$ is the capacity of each $\epsilon_{\textnormal{conv-quota}}$ filter. It follows that the number of conversion sites with non-zero privacy loss is precisely $N^{\leq t,\mathsf{adv}}$, so:
(93) $\leq\left|\left\{c\in\mathbf{bad_{c}}:\epsilon_{\textnormal{conv-quota}}[c]^{\leq T-1}>0\right\}\right|\cdot\epsilon_{\textnormal{conv-quota}}$
(94) $=N^{\leq T-1,\mathsf{adv}}\cdot\epsilon_{\textnormal{conv-quota}}.$
Now, during the 2-PC for time $T$, we have the following cases:
  • Suppose $\epsilon^{T}_{x}$ is a reasonable value, in the sense that it’s bounded by the capacity $\epsilon_{\textnormal{conv-quota}}$. Then,
    (95) $\epsilon^{\leq T,\mathbf{bad}}_{\mathsf{global}}=\epsilon_{\text{used}}^{\mathbf{bad}}+\epsilon^{T}_{x}\leq N^{\leq T,\mathsf{adv}}\cdot\epsilon_{\textnormal{conv-quota}}.$
  • Otherwise, $\epsilon^{T}_{x}$ is unreasonable, in which case $\epsilon^{T}_{x}$ exceeds the capacity $\epsilon_{\textnormal{conv-quota}}$. In this case,
    (96) $\epsilon_{\textnormal{conv-quota}}[c_{t}]^{\leq T-1}+\epsilon^{T}_{x}\geq\epsilon_{\textnormal{conv-quota}},$
    causing $\mathcal{F}^{\epsilon_{\textnormal{conv-quota}}[c]}_{x}.\operatorname{canConsume}(\epsilon^{T}_{x})$ to return FALSE by definition, so no budget is spent at all. In such a case,
    (97) $\epsilon^{\leq T,\mathbf{bad}}_{\mathsf{global}}=\epsilon_{\text{used}}^{\mathbf{bad}}+0=\epsilon_{\text{used}}^{\mathbf{bad}}\leq N^{\leq T-1,\mathsf{adv}}\cdot\epsilon_{\textnormal{conv-quota}},$
    by equation (94).
Since the adversary has created at most $N^{\mathsf{adv}}$ by the end of time $T$, it must be the case that $N^{\leq T-1,\mathsf{adv}}\leq N^{\leq T,\mathsf{adv}}\leq N^{\mathsf{adv}}$. This means that, in either case, the attackers can consume at most $N^{\leq T,\mathsf{adv}}\epsilon_{\textnormal{conv-quota}}\leq N^{\mathsf{adv}}\epsilon_{\textnormal{conv-quota}}$ of the global budget budget by the end of time $T$, as desired ∎
Lemma 0 (Impression-Site Quota Allocation Consistency).
Fix a record $x=(d,e,F)$ and a report $\rho$ at step $k$. Denote by $\epsilon_{x}^{k}$ the epoch-level privacy loss given by EpochBudget, denote by $\epsilon_{x}^{\mathbf{i},k}[i]$ the epoch-impression-site-level privacy loss given by EpochImpSiteBudget (Def. 1) using $2\Delta(\rho)$ as an upper bound for $\Delta_{i}(\rho)$ (Lem. 10). We have:
(98) $\epsilon_{x}^{k}\leq\sum_{i\in\mathcal{S}}\epsilon_{x}^{\mathbf{i},k}[i]$
Proof.
We go through the different cases for the upper bound on $\Delta_{x}(\rho)$:
  • If $d\neq d_{r},e\neq E_{r}$ or $F\cap F_{A}=\emptyset$, then $\epsilon_{x}^{k}=0$. In that case, by Thm. 9 we also have $\epsilon_{x}^{\mathbf{i},k}[i]=0$ for all $i$.
  • If $d=d_{r}$ and $E_{r}=\{e\}$:
    • If $I_{r}=\{i\}$, we have $\epsilon_{x}^{k}=\epsilon_{x}^{\mathbf{i},k}[i]=\|A(F)-A(\emptyset)\|_{1}/\lambda$.
    • Else, we have $\epsilon_{x}^{\mathbf{i},k}[i]=2\Delta(\rho)/\lambda\geq\|A(F)-A(\emptyset)\|_{1}/\lambda=\epsilon_{x}^{k}$.
  • Else, we have $\epsilon_{x}^{\mathbf{i},k}[i]=2\Delta(\rho)/\lambda\geq\Delta(\rho)/\lambda=\epsilon_{x}^{k}$.
Remark that the upper bound in Lem. 6 can be quite loose. Since the impression-site quota has no privacy meaning and is only used through Lem. 6 to obtain isolation guarantees, instead of using per-epoch-site privacy loss we could use any heuristic that also satisfies Lem. 6. For instance, we could define $\epsilon_{x}^{\mathbf{i},k}[i]$ by dividing $\epsilon_{x}^{k}$ uniformly across impression sites $i$ with non-zero contributions.
Lemma 0.
Consider a sequence of T batch query steps. If an adversary, across all individual report generation attempts within these T steps, successfully leads to the instantiation (via $\operatorname{saveImpression}$ calls) and subsequent use (successful budget consumption from) of at most $M^{\text{adv}}$ distinct imp-quota filters and $N^{\text{adv}}$ distinct conv-quota filters in an attempt to deplete the global budget, then the adversary consumes at most $N^{\mathsf{adv}}\epsilon_{\textnormal{conv-quota}}$ budget from the global budget.
Proof.
By basic composition under a pure DP filter, $\epsilon_{\textnormal{global budget}}^{\leq t-1,\mathsf{bad}}$ is the sum of global budget consumption by reports associated with adversarial conversion sites, $\mathsf{bad_{c}}$ across all query steps $s$ from $1$ up to t1t-1. Let $R_{s}$ be the set of reports in query step $s$, $c_{r,s}$ be the conversion site for report $r$ from query step $s$, $\epsilon_{r,s}^{\textnormal{global budget}}$ be the global budget consumed by that report, and $\text{pass}(r,s)$ indicate if it was successfully processed. Then,
(99) $\epsilon_{\mathsf{global}}^{\leq t-1,\mathbf{bad}}$ $=\sum_{s\in[t-1]}\sum_{r\in R_{s}:c_{r,s}\in\mathsf{bad_{c}}}\epsilon_{r,s}^{\textnormal{global budget}}\cdot\text{pass}(r,s)$
(100) $\leq\sum_{s\in[t-1]}\sum_{i\in I_{x}^{s}\cap\mathbf{bad_{i}}}\epsilon_{\mathbf{x}}^{\mathbf{i},s}[i]\cdot\text{pass}(s),$
where the last inequality follows from Lem. 6. First, by the restriction in the sum, we know $s$ satisfies $c_{s}\in\mathbf{bad_{c}}$. Second, recall that, for a conversion site to incur privacy losses on impression sites, the conversion site must register these impression sites, meaning that if $c_{s}\in\mathbf{bad_{c}}$, then $\mathbf{i}_{s}\subseteq\mathbf{bad_{i}}$. Now, continuing where we ended in Eq. 100, we get:
(101) $\epsilon_{\mathsf{global}}^{\leq t-1,\mathbf{bad}}$ $\leq\sum_{s\in[t-1]:c_{s}\in\mathbf{bad_{c}}}\sum_{i\in\mathbf{bad_{i}}}\epsilon_{\mathbf{x}}^{\mathbf{i},s}[i]\cdot\text{pass}(s)$
(102) $=\sum_{i\in\mathbf{bad_{i}}}\sum_{s\in[t-1]:c_{s}\in\mathbf{bad_{c}},i\in\mathbf{i}_{s}}\epsilon_{\mathbf{x}}^{\mathbf{i},s}[i]\cdot\text{pass}(s),$
(103) $\leq\sum_{i\in\mathbf{bad_{i}}}\sum_{s\in[t-1]:i\in\mathbf{i}_{s}}\epsilon_{\mathbf{x}}^{\mathbf{i},s}[i]\cdot\text{pass}(s),$
by changing order of summation, and the last inequality by relaxing the $c_{s}\in\mathbf{bad_{c}}$” condition. But note that
(104) $\epsilon_{\textnormal{imp-quota}}^{\leq t-1}[i]=\sum_{s\in[t-1]:i\in\mathbf{i}_{s}}\epsilon_{\mathbf{x}}^{\mathbf{i},s}[i]\cdot\text{pass}(s),$
because, by epoch-site-level consistency property in lemma 4, we know that only relevant site $i$ at time $s$, where every filter has enough budget to pass the 2-PC check, will have epoch-site level privacy losses incurred. Substituting this equality into equation (103), we get:
(105) $\epsilon_{\mathsf{global}}^{\leq t-1,\mathbf{bad}}$ $\leq\sum_{i\in\mathbf{bad_{i}}}\epsilon_{\textnormal{imp-quota}}^{\leq t-1}[i]$
(106) $=\sum_{i\in\mathbf{bad_{i}}:\epsilon_{\textnormal{imp-quota}}^{\leq t-1}[i]>0}\epsilon_{\textnormal{imp-quota}}^{\leq t-1}[i]$
(107) $\leq\sum_{i\in\mathbf{bad_{i}}:\epsilon_{\textnormal{imp-quota}}^{\leq t-1}[i]>0}\epsilon_{\textnormal{imp-quota}}$
(108) $=\left|\left\{i\in\mathbf{bad_{i}}:\epsilon_{\textnormal{imp-quota}}^{\leq t-1}[i]>0\right\}\right|\cdot\epsilon_{\textnormal{imp-quota}},$
because only non-zero privacy losses that were incurred contribute meaningfully to the composition. Finally, we note that $\left|\left\{i\in\mathbf{bad_{i}}:\epsilon_{\textnormal{imp-quota}}^{\leq t-1}[i]>0\right\}\right|\leq M^{\leq t-1,\mathsf{adv}}$ by definition and:
(109) $\epsilon_{\mathsf{global}}^{\leq t-1,\mathbf{bad}}$ $\leq M^{\leq t-1,\mathsf{adv}}\cdot\epsilon_{\textnormal{imp-quota}}.$
Following this result, similar to the proof for part 1:
  • Suppose $\epsilon^{t}_{x}\leq\epsilon_{\textnormal{imp-quota}}$,
    (110) $\epsilon^{\leq t,\mathbf{bad}}_{\mathsf{global}}\leq M^{\leq t,\mathsf{adv}}\cdot\epsilon_{\textnormal{imp-quota}}.$
  • Else, $\epsilon^{t}_{x}>\epsilon_{\textnormal{imp-quota}}$, then $\epsilon_{\textnormal{imp-quota}}$ will be exceeded, causing $\operatorname{canConsume}$ to return FALSE, so,
    (111) $\epsilon_{\mathsf{global}}^{\leq t,\mathbf{bad}}=\epsilon^{\leq t-1,\mathbf{bad}}_{\mathsf{global}}+0=\epsilon^{\leq t-1,\mathbf{bad}}_{\mathsf{global}}\leq M^{\leq t-1,\mathsf{adv}}\cdot\epsilon_{\textnormal{imp-quota}},$
    by equation (109).
Since by the end of time $t$, the adversary has created at most $M^{\mathsf{adv}}$ imp-quota filters, we know $M^{\leq t-1,\mathsf{adv}}\leq M^{\leq t,\mathsf{adv}}\leq M^{\mathsf{adv}}$, which means that in both cases we have:
(112) $\epsilon^{\leq t,\mathbf{bad}}_{\mathsf{global}}\leq M^{\leq t,\mathsf{adv}}\epsilon_{\textnormal{imp-quota}}\leq M^{\mathsf{adv}}\epsilon_{\textnormal{imp-quota}},$
as desired. ∎
We can now combine Lem. 5 and 7 to obtain our main isolation theorem:
Theorem 1 (Resilience to DoS depletion).
Consider an adversary who manages to create $M^{\textrm{adv}}$ and $N^{\textrm{adv}}$ imp-quota and conv-quota filters, respectively. The maximum budget $\epsilon_{\textnormal{global}}^{\textrm{adv}}$ that the adversary can consume from the global budget on a device $d$ is such that:
(113) $\epsilon_{\textnormal{global}}^{\textrm{adv}}\leq\min(M^{\textrm{adv}}\epsilon_{\textnormal{imp-quota}},\ N^{\textrm{adv}}\epsilon_{\textnormal{conv-quota}}).$
Proof.
At any time $t$, if the adversary controls at most $M^{\leq t,\mathsf{adv}}\leq M^{\mathsf{adv}}$ imp-quota and $N^{\leq t,\mathsf{adv}}\leq N^{\mathsf{adv}}$ conv-quota filters, then by Lem. 5 and 7:
(114) $\epsilon^{\mathsf{adv}}_{\mathsf{global}}$ $\leq M^{\leq t,\mathsf{adv}}\epsilon_{\textnormal{imp-quota}}\leq M^{\mathsf{adv}}\epsilon_{\textnormal{imp-quota}}$
(115) $\epsilon^{\mathsf{adv}}_{\mathsf{global}}$ $\leq N^{\leq t,\mathsf{adv}}\epsilon_{\textnormal{conv-quota}}\leq N^{\mathsf{adv}}\epsilon_{\textnormal{conv-quota}}.$
Therefore:
(116) $\epsilon_{\textnormal{global}}^{\textrm{adv}}\leq\min(M^{\textrm{adv}}\epsilon_{\textnormal{imp-quota}},\ N^{\textrm{adv}}\epsilon_{\textnormal{conv-quota}}).$
Appendix D Cross-report Privacy Loss Optimization
In §C, we presented a general algorithm where queriers pay for each report separately. This section formalizes the cross-report optimization sketched in §4.5, by defining variations of Alg. 8, Alg. 6 in §D.1. We focus on histogram reports, defined in D.2, and show that paying only once for a sequence of properly correlated histogram reports still satisfies global IDP guarantees in §D.3. We leave the generalization to other queries for future work.
D.1  Algorithm
Overview. Alg. 9 updates the formalism from Alg. 8. Instead of generating a report immediately upon conversion, the first querier to request a report attached to a report identifier $r$ calls MeasureConversion to create a stateful attribution object $\alpha_{r}$, which pays upfront for any valid sequence of histogram reports over non-overlapping impressions.
MeasureConversion performs attribution using the pre-specified attribution function $a$ and the noise scale $\lambda$, which give an upper bound on the total leakage of any sequence of future reports applying $a$ on disjoint sets of impressions. We run through a modified two-phase commit protocol, to deduct budget from the global filter and the quotas, but not from the querier filters.  [^2] The querier budget is left out of the 2PC because we don’t know ahead of time which queriers will request a report, and we don’t want to block some queriers if other queriers are out of budget. Finally, $\alpha_{r}$ stores the attribution function $a_{\mathbf{F}}$ with corresponding impressions $\mathbf{F}$, the privacy parameters $\lambda$, and the support $S$ of impressions requested so far.
GetReport allows a querier to receive a report from $r$, once the attribution object $\alpha_{r}$ has been created. $\alpha_{r}$ is only created once, and is reused every time a querier requests a report from $r$. GetReport checks that privacy parameters match what was committed in the attribution object, and that impressions are not queried twice. If these checks pass, GetReport spends querier budget and computes the report using the predefined attribution function $a$ stored in $\alpha_{r}$. We update the support of impressions $S\leftarrow S\sqcup S_{\rho}$ in $\alpha_{r}$ each time a new report $\rho$ requests impressions in $S_{\rho}\subset\mathcal{I}$
Subroutine. We define AtomicFilterCheckAndConsume2 as in Alg. 7, except that querier filters are not part of the atomic commit. That is, we do not take $b$ as an input, and we delete Line 6 and Line 16 from Alg. 7.
Alg. 9 Updated formalism for cross-report optimization (diff with Alg. 8)
1:// Collect, aggregate and noise reports to answer $Q$
2:def $\operatorname{AnswerQuery}$($D,Q,\lambda,b$)
[…]
4:for $r\in R$ do
5:if $\alpha_{r}$ is not defined then
6:  // Initialize attribution object
7:  $\alpha_{r}\leftarrow\operatorname{MeasureConversion}(D,\rho_{r}^{b},\lambda)$
[…]
9: Sample $X\sim\mathcal{L}(\lambda)$
10:return $\sum_{r\in R}\rho_{r}^{b}(D)+X$
Alg. 10 Big Bird algorithm (on-device) with cross-report optimization for histogram reports
1:Input
2: Filter and quota capacities $\epsilon_{\textnormal{global}}$, $\epsilon_{\textnormal{querier}}$, $\epsilon_{\textnormal{imp-quota}}$, $\epsilon_{\textnormal{conv-quota}}$
3:def $\operatorname{MeasureConversion}$($D,\rho,\lambda,\mathbf{b}$)
4: Read $\rho$ to get device $d$, epoch $E$, conversion site $c$, impression sites $\mathbf{i}$, histogram attribution function $A_{a_{\rho},S_{\rho},H_{\rho}}$ with scalar attribution function $a$, support impressions $S$ and histogram bin mapping $H$.
5:for $e\in E$ do
[…]
7:if $\mathcal{F}_{x}$ is not defined then
8:  $\mathcal{F}_{x}\leftarrow\operatorname{InitializeFilters(\epsilon_{\textnormal{global}},\epsilon_{\textnormal{querier}},\epsilon_{\textnormal{imp-quota}},\epsilon_{\textnormal{conv-quota}})}$
[…]
11:if $\operatorname{AtomicFilterCheckAndConsume2}(\mathcal{F}_{x},c,\mathbf{i},\epsilon_{x},\epsilon_{\mathbf{x}}^{\mathbf{i}})=\textsf{FALSE}$ then
12:  $F_{e}\leftarrow\emptyset$ // Empty the epoch if any filter check fails
13:return $\alpha=(a_{\mathbf{F}},\mathbf{F},\lambda,\emptyset)$ // Start with $S_{\alpha}=\emptyset$
14:
15:def $\operatorname{GetReport}$($\alpha,\rho,\lambda$)
16: Read $\alpha$ to get attribution function $a_{\alpha}$, impressions $\mathbf{F}_{\alpha}$, noise scale $\lambda_{\alpha}$.
17: Read $\rho$ to get device $d$, querier $b$, target epochs $E$, histogram attribution function $A_{a,S,H}$ with scalar attribution function $a$, support impressions $S$ and histogram bin mapping $H$.
18:if $S\cap S_{\alpha}\neq\emptyset\vee\lambda\neq\lambda_{\alpha}\vee a_{\rho}\neq a_{\alpha}$ then
[…]
20:return $\rho$
21:for $e\in E$ do
[…]
24:if $\mathcal{F}_{x}^{\textnormal{querier}[b]}.\operatorname{tryConsume}(\epsilon^{t}_{x})=\textsf{FALSE}$ then
25:  $F_{e}\leftarrow\emptyset$
26:else
27:  $F_{e}\leftarrow{\mathbf{F}_{\alpha}}_{e}$
[…]
30:return $\rho,\alpha$
D.2  Histogram reports definition and properties
The following definitions (Def. 1 and 2) are adapted from (TKM+24, 41, Thm. 18). Histogram reports distribute a positive value across impressions, map each impression to a bucket, and then sum up the attributed value in each bucket. Lem. 3 gives the global sensitivity of such histogram reports, which is bounded by the maximum attributable value.
Definition 0 (Scalar attribution function).
Fix $k>0$ a number of epochs. A scalar attribution function is a function $a:\mathcal{P}(\mathcal{I})^{k}\times\mathcal{I}\rightarrow\mathbb{R}_{+}$ that attributes a positive value $a_{\mathbf{F}}(f)$ to each impression $f\in\mathcal{I}$, depending on all the impressions in $k$ epochs $\mathbf{F}\in\mathcal{P}(\mathcal{I})^{k}$.
For a scalar attribution function $a$, we define its maximum attributable value $a^{\max}$ as follows:
(117) $a^{\max}:=\max_{\mathbf{F}\in\mathcal{P}(\mathcal{I})^{k}}\sum_{j=1}^{k}\sum_{f\in\mathbf{F}_{j}}a_{\mathbf{F}}(f)$
Definition 0 (Histogram report).
[…]
First, we define $A_{a,S,H}:\mathcal{P}(\mathcal{I})^{k}\to\mathbb{R}^{m}$ as follows:
(118) $A(\mathbf{F})=\sum_{j=1}^{k}\sum_{f\in\mathbf{F}_{j}}\mathds{1}[f\in S]a_{\mathbf{F}}(f)\cdot H(f)$
$A_{a,S,H}$ is a well-defined attribution function (in the sense of Def. 2).
Second, for a device $d$ and a set of epochs $E$ we define the histogram report associated with $A_{a,S,H}$, as in Def. 3:
(119) $\rho:D\mapsto A_{a,S,H}(D_{d}^{E})$
Next, Lem. 3 and 4 give two preliminary properties of histogram reports, that will be used in Thm. 5.
Lemma 0 (Histogram sensitivity).
Consider a histogram report $\rho$ with associated attribution function $A_{a,S,H}$. We have:
(120) $\Delta(\rho)\leq 2a^{\max}$
Proof.
Take a report $\rho$ with scalar attribution function $a$, device $d$ and epochs $E$. Consider two neighborhing databases $D,D^{\prime}$ and denote $\mathbf{F}:=D_{d}^{E}$ and $\mathbf{F}^{\prime}:={D^{\prime}}_{d}^{E}$. We have:
(121) $\|\rho(D)-\rho(D^{\prime})\|_{1}=\|A_{a,S,H}(\mathbf{F})-A_{a,S,H}(\mathbf{F}^{\prime})\|_{1}$
(122) $=\left\|\sum_{j=1}^{k}\sum_{f\in\mathbf{F}_{j}}a_{\mathbf{F}}(f)\cdot H(f)-\sum_{j=1}^{k}\sum_{f\in\mathbf{F}^{\prime}_{j}}a_{\mathbf{F}^{\prime}}(f)\cdot H(f)\right\|_{1}$
(123) $\leq\sum_{j=1}^{k}\sum_{f\in\mathbf{F}_{j}}a_{\mathbf{F}}(f)\|H(f)\|+\sum_{j=1}^{k}\sum_{f\in\mathbf{F}^{\prime}_{j}}a_{\mathbf{F}^{\prime}}(f)\|H(f)\|$
(124) $\leq 2a^{\max}$
Even though this bound holds even for non-neighborhing databases, (TKM+24, 41, Thm. 18) provides mild conditions under which the bound is tight. ∎
D.3  Privacy proof
Lemma 0 (Correlated histogram sensitivity).
Fix a device-epoch $x=(d,e,F)$ and a database $D$. Fix a report identifier $r\in\mathbb{Z}$ corresponding to a histogram attribution object $\alpha_{r}$. Fix a sequence of reports $\rho_{1},\dots,\rho_{n}$ that request a report from $r$, ordered by lexicographically by time and querier $(t,b)$. In particular, MeasureConversion is called for $\rho_{1}$ and then reused for subsequent reports. $\sum_{i=1}^{n}\|\rho_{i}(D)-\rho_{i}(D+x)\|/\lambda_{i}$ represents the total contribution over reports computed from $\alpha_{r}$, each with its own requested noise scale $\lambda_{i}$. Denote by $\text{pass}_{1}$ the output of the 2PC for $x$ in Alg. 10.
We have:
$\sum_{i=1}^{n}$ $\|\rho_{i}(D)-\rho_{i}(D+x)\|/\lambda_{i}$
(125) $\leq\text{pass}_{1}\cdot 2a^{\max}/\lambda_{1}$
Proof.
If $\text{pass}_{1}=0$, then $\rho_{i}(D)=\rho_{i}(D+x)$ because in both cases the data for $x$ is zeroed-out ($F_{e}=\emptyset$ at Line 12), and we’re done.
Now, suppose that $\text{pass}_{1}=1$. Take a report $\rho_{i}$ with scalar attribution function $a_{i}$ and support impressions $S_{i}$. These only depend on past results $v_{<t_{i},b_{i}}$. If $S_{i}\cap(S_{1}\cup\dots\cup S_{i-1})\neq\emptyset$, $\lambda_{i}\neq\lambda_{1}$ or $a_{i}\neq a_{1}$, then $\rho_{i}(D)=A(\emptyset,\dots,\emptyset)=\rho_{i}(D+x)$. Also, if $x=(d,e,F)$ is not queried by $\rho_{i}$ ($d\neq d_{i}$ or $e\not\in E_{i}$), then $\rho_{i}(D)=\rho_{i}(D+x)$.
Denote by $I$ the set of remaining reports verifying $S_{i}\cap(S_{1}\sqcup\dots S_{i-1})=\emptyset$, $\lambda_{i}=\lambda_{1}$, $a_{i}\neq a_{1}$, $d_{i}=d$ and $e\in E_{i}$. We have:
(126) $\sum_{i=1}^{n}\|\rho_{i}(D)-\rho_{i}(D+x)\|/\lambda_{i}$
(127) $=\sum_{i\in I}\|\rho_{i}(D)-\rho_{i}(D+x)\|/\lambda_{1}$
(128) $=\sum_{i\in I}\|\sum_{j=1}^{k}\sum_{f\in\mathbf{F}_{j}}\mathds{1}[f\in S_{i}]a_{\mathbf{F}}(f)H_{i}(f)$-
(129) $\sum_{j=1}^{k}\sum_{f\in\mathbf{F}_{j}^{\prime}}\mathds{1}[f\in S_{i}]a_{\mathbf{F}^{\prime}}(f)H_{i}(f)\|/\lambda_{1}$
(130) $\leq\sum_{i\in I}\sum_{j=1}^{k}\sum_{f\in\mathbf{F}_{j}}\mathds{1}[f\in S_{i}](a_{\mathbf{F}}(f)+a_{\mathbf{F}^{\prime}}(f))/\lambda_{1}$
(131) $\leq\sum_{j=1}^{k}\sum_{f\in\mathbf{F}_{j}}\mathds{1}[f\in\sqcup_{i\in I}S_{i}](a_{\mathbf{F}}(f)+a_{\mathbf{F}^{\prime}}(f))/\lambda_{1}$
(132) $\leq 2a^{\max}/\lambda_{1}$
by Def. 1 and using the fact that the $S_{i}$ are disjoint so each impression is counted at most once.
Theorem 5.
Consider $x\in\mathcal{X}$ with global budget capacity $\epsilon_{\textnormal{global}}$. For simplicity, since we are considering global guarantees, we consider a non-adaptive data generation process in Alg. 9, which defines a mechanism $q$. Then, $q$ satisfies individual device-epoch $\epsilon_{\textnormal{global}}$-DP for $x$ under public information $\mathcal{C}$.
Proof.
Take a device-epoch $x=(d,e,F)\in\mathcal{X}$ and a database $D$ that doesn’t contain $(d,e)$. Denote by $x_{\mathcal{C}}=(d,e,F\cap\mathcal{C})$ the device-epoch obtained by keeping only public events $\mathcal{C}$ from $x$, where public events are the set of all conversions. Take $v\in\operatorname{Range}(q)$. As in Thm. 3, want to show that:
(133) $\left|\ln\left(\frac{\Pr[q(D+x_{\mathcal{C}})=v]}{\Pr[q(D+x)=v]}\right)\right|\leq\epsilon_{\textnormal{global}}.$
Using Bayes’ rule and $\rho(D+x_{\mathcal{C}})=\rho(D)$, as in Thm. 3, we have:
(134) $\left|\ln\left(\frac{\Pr[q(D+x_{\mathcal{C}})=v]}{\Pr[q(D+x)=v]}\right)\right|$
(135) $=$ $\left|\ln\left(\frac{\prod_{t=1}^{t_{\max}}\prod_{b\in S}\Pr[\sum_{r\in R_{t}^{b}}\rho_{r,t}^{b}(D)+X_{t}^{b}=v_{t}^{b}]}{\prod_{t=1}^{t_{\max}}\prod_{b\in S}\Pr[\sum_{r\in R_{t}^{b}}\rho_{r,t}^{b}(D+x)+X_{t}^{b}=v_{t}^{b}]}\right)\right|$
(136) $\leq$ $\sum_{t=1}^{t_{\max}}\sum_{b\in S}\left|\ln\left(\frac{\Pr[\sum_{r\in R_{t}^{b}}\rho_{r,t}^{b}(D)+X_{t}^{b}=v_{t}^{b}]}{\Pr[\sum_{r\in R_{t}^{b}}\rho_{r,t}^{b}(D+x)+X_{t}^{b}=v_{t}^{b}]}\right)\right|$
where each query $Q_{t}^{b}=\{\rho_{r,t}^{b},r\in R_{t}^{b}\}$ is chosen adaptively, potentially based on previous results $v_{1}^{b},\dots,v_{t-1}^{b}$. Since the filters and attribution functions are identical for $D$ and $D+x$ when we condition on past results, we write $\rho_{r,t}^{b}(D)$ for simplicity instead of $\rho_{r,t}^{b}(D;\mathcal{F}_{v_{<t}},\alpha_{v_{<t}}))$.
Fix $t\in[t_{\max}]$ and $b\in\mathcal{S}$. Without the optimization, as in Eq. 80, the device would pay for each report sent to any querier, which would give:
(137) $\left|\ln\left(\frac{\Pr[\sum_{r\in R_{t}^{b}}\rho_{r,t}^{b}(D)+X_{t}^{b}=v_{t}^{b}]}{\Pr[\sum_{r\in R_{t}^{b}}\rho_{r,t}^{b}(D+x)+X_{t}^{b}=v_{t}^{b}]}\right)\right|$
(138) $\leq\sum_{r\in R_{t}^{b}}\Delta_{x}(\rho_{r,t}^{b})\text{pass}_{r}^{b}/\lambda_{t}^{b}$
Instead of upper-bounding the difference $\|\rho_{r,t}^{b}(D)-\rho_{r,t}^{b}(D+x)\|$ for each report by $\Delta_{x}(\rho_{r})$ separately, which takes a maximum over all $D$ right away, we keep information about $D$ a bit longer. This will allow us to leverage the fact that reports $\rho_{r,t}^{b}$ across different timesteps and queriers tied to a same identifier $r$ are correlated:
(139) $\left|\ln\left(\frac{\Pr[\sum_{r\in R_{t}^{b}}\rho_{r,t}^{b}(D)+X_{t}^{b}=v_{t}^{b}]}{\Pr[\sum_{r\in R_{t}^{b}}\rho_{r,t}^{b}(D+x)+X_{t}^{b}=v_{t}^{b}]}\right)\right|$
(140) $\leq\sum_{r\in R_{t}^{b}}\|\rho_{r}^{b}(D)-\rho_{r}^{b}(D+x)\|/\lambda_{t}^{b}$
For a report identifier $r\in\mathbb{Z}$, we now define $\mathcal{T}_{r}$, which keeps track of all queriers that requested a report from $r$ and at which timesteps they requested it:
(141) $\mathcal{T}_{r}:=\{(t,b)\in[t_{\max}]\times\mathcal{S}:r\in R_{t}^{b}\}$
This notation allows us to swap the sums, after putting Eq. 140 into Eq. 136:
(142) $\left|\ln\left(\frac{\Pr[q(D+x_{\mathcal{C}})=v]}{\Pr[q(D+x)=v]}\right)\right|$
(143) $\leq\sum_{t=1}^{t_{\max}}\sum_{b\in S}\sum_{r\in R_{t}^{b}}\|\rho_{r}^{b}(D)-\rho_{r}^{b}(D+x)\|/\lambda_{t}^{b}$
(144) $=\sum_{r\in\mathbb{Z}}\sum_{(t,b)\in\mathcal{T}_{r}}\|\rho_{r}^{b}(D)-\rho_{r}^{b}(D+x)\|/\lambda_{t}^{b}$
Fix a report identifier $r\in\mathbb{Z}$ corresponding to a histogram attribution object $\alpha_{r}$. Denote by $t_{0},b_{0}$ the first time step and first querier that requests a report from $r$, thereby calling MeasureConversion. By Lem. 4, we have:
(145) $\sum_{(t,b)\in\mathcal{T}_{r}}\|\rho_{r}^{b}(D)-\rho_{r}^{b}(D+x)\|/\lambda_{t}^{b}\leq\text{pass}_{r,t_{0}}^{b_{0}}\cdot\Delta_{x}(\alpha_{r})/\lambda_{t_{0}}^{b_{0}}$
Finally, since $\text{pass}_{r,t_{0}}^{b_{0}}$ implies that $\epsilon_{r,t_{0}}^{b_{0}}=\Delta_{x}(\alpha_{r})/\lambda_{t_{0}}^{b_{0}}$ passes the global budget, by definition of the global budget Eq. 144 becomes:
(146) $\left|\ln\left(\frac{\Pr[q(D+x_{\mathcal{C}})=v]}{\Pr[q(D+x)=v]}\right)\right|$
(147) $\leq\sum_{r\in\mathbb{Z}}\text{pass}_{r,t_{0}}^{b_{0}}\epsilon_{r,t_{0}}^{b_{0}}$
(148) $\leq\epsilon_{\textnormal{global}}$
which concludes the proof. ∎
Remark. Lem. 3 and 4 show that for histogram reports with $k>1$ epochs, Alg. 10 spends up to $|\mathbf{b}|$ times less budget than Alg. 6 when $|\mathbf{b}|$ queriers request reports from the same conversion with a single report identifier $r$. This is because the privacy loss in these cases is proportional to $\Delta(\rho)=2a^{\max}$. For histogram reports with a single epoch, we can use the individual sensitivity, which renders this optimization unnecessary.
Appendix E Prototype Screeenshot
Fig. 6. Firefox privacy loss dashboard.
Fig. 6 shows a screenshot of our Firefox extension that serves as a dashboard for visualizing privacy loss across the different filters and quotas Big Bird maintains. The screenshot follows a scenario of user visits and purchases, which we emulate programmatically on our local browsers, since no site currently invokes the PPA API. The scenario is as follows: A user visits many websites that display ads on them, such as nytimes.com and blog.com. These websites store every ad view as an event using saveImpression(). The user then purchases products for which they have seen ads, including on nike.com and toys.com. At time of purchase, these websites call measureConversion() to generate and send a report, consuming privacy in the process. The user wants to check how much of their privacy budget has been spent using the dashboard in Fig. 6.